### v1

# Article Type: Pre-Print

Abstract
Cybersecurity threats in automotive systems pose significant risks to safety and reliability. This article introduces a methodology integrating threat-informed dynamic security modelling with a Threat Analysis and Risk Assessment workflow. Using the example of an In-Vehicle Infotainment system, we demonstrate the methodology’s application in risk management to strengthen automotive resiliency.
Keywords: risk management, security, autonomous vehicles
journal: IEEE Security & Privacy
Introduction
Automotive systems have transformed into sophisticated platforms brimming with cutting-edge features. In contrast to their predecessors, today’s automotive systems are equipped with a variety of advanced sensors and connectivity options, such as Control Area Networks (CAN), Local Interconnect Networks (LIN), FlexRay, and Ethernet. They also offer seamless external connectivity through Bluetooth, Wi-Fi, and GPS. Additionally, the integration of feature-rich In-Vehicle Infotainment (IVI) systems and powerful computing units for autonomous driving continues to push the boundaries of what is possible in automotive technology. However, this technological evolution also broadens the attack surface, making automotive systems more susceptible to cybersecurity attacks.
Cybersecurity threats and vulnerabilities in the automotive industry are well-documented, and attackers exploit these vulnerabilities to compromise vehicle security, access sensitive data, and even take control of vehicle functions.
Recent literature has increasingly highlighted the challenges of defending against cybersecurity threats in automotive systems. In their survey on security attacks and defence techniques for Connected and Autonomous Vehicles (CAVs), Pham and Xiong (2021) observed that attackers could potentially compromise one vehicle component after another. To mitigate this risk, they emphasised the need for manufacturers to strengthen the security of all system components to ensure the overall protection of CAVs.
While this approach may be feasible during the early stages of an automotive system’s life cycle, evolving cybersecurity threats are likely to render it less effective over time. Given the rapid increase in both threats and the complexity of automotive systems, it is unrealistic to expect that an automotive system’s initial cybersecurity features will counter all emerging risks throughout its entire life cycle. This is where threat-informed dynamic threat and attack models become crucial, offering a proactive and adaptive approach to designing secure automotive systems.
Dynamic threat and attack models provide engineers with a structured framework to analyse and mitigate cybersecurity risks, ensuring that automotive systems remain protected against evolving cybersecurity threats. Additionally, they enable consumers to make informed decisions when selecting vehicles based on their cybersecurity robustness.
However, many of the threat and attack models discussed in the literature do not fully address the unique cybersecurity requirements of modern automotive systems. In our previous work [1], we conducted a systematic literature review of threat and attack modelling approaches pertinent to cyber-physical systems (CPSs)—systems that integrate computational and physical process—of which automotive systems are an example. Our investigation yielded several findings, including:
  • Threat modelling is commonly conducted only in early stages of system development, and does not always consider attack models.
  • Threat and attack models in the literature primarily focus on IT systems, making them challenging to apply to automotive systems.
  • Much of the existing literature does not differentiate between cybersecurity breaches and consequences in IT systems versus those in automotive systems.
  • Ambiguity persists in the literature regarding the definitions and relationships between threat modelling and attack modelling.
These findings highlight the need to tailor security modeling approaches for automotive systems so that they can adequately capture the unique cybersecurity threats faced in these environments.
For instance, threat models developed using Microsoft’s STRIDE (Spoofing, Tampering, Repudiation, Information Disclosure, Denial of Service, and Elevation of Privilege) primarily focus on data flows and information security, overlooking critical aspects of how such breaches can directly affect automotive functions and safety. In other words, threat modelling of automotive systems often do not consider consequences of cybersecurity attacks beyond information security. On a similar note, attack models designed with traditional IT systems in mind do not fully align with multi-path tactics and techniques employed by attackers seeking to gain control of a vehicle. These techniques often involve a combination of physical and cyber actions.
In short, our investigation has highlighted limitations in current threat and attack modelling methods that underscore the need for a tailored, dynamic approach. Threat and attack models need to iteratively and continuously consider both consequences and characteristics of cyber-physical attacks so as to offer a comprehensive approach to safeguarding automotive systems.
Contributions. In this article, we introduce ACTISM (Automotive Consequence-Driven and Threat-Informed Security Modelling) — a novel security modeling approach that enhances the resilience of automotive systems by dynamically updating their cybersecurity posture in response to prevailing threats, attacker tactics, and their impact on system functionality and safety. We demonstrate the effectiveness of ACTISM by applying it to an IVI system, illustrating how the model can be adapted as new threats emerge.
Unlike existing security modelling approaches, ACTISM dynamically updates security models with new information acquired through threat intelligence and ongoing security monitoring. This keeps security models relevant against evolving attacker tactics, techniques and procedures (TTPs) even if the models were developed in early stages of system development. ACTISM also allows practitioners to apply the security models in both IT and automotive domains by using attack paths to capture the multi-path and multi-agent aspects of real-world cyber-physical attacks.
Cybersecurity incidents in automotive systems differ significantly from those in traditional IT systems. These incidents can lead to complex failure modes with both cyber and physical consequences, such as vehicle misnavigation and potentially dangerous outcomes. By adopting a consequence-driven and cyber-informed approach to cybersecurity, ACTISM ensures that cyber-physical attacks, effects and consequences are considered in security modelling for automotive systems.
ACTISM seamlessly and systematically integrates with, and adapts to a number of threat analysis and risk assessment frameworks such as:
  • Microsoft’s STRIDE methodology.
  • HEAVENS (HEAling Vulnerabilities to ENhance Software Security and Safety), a structured and model-based hazard analysis and risk assessment method for automotive systems.
  • Attack Tree, a hierarchical diagram representing potential attack paths, illustrating how an attacker could exploit system vulnerabilities to achieve specific goals.
  • Common Vulnerability Scoring System (CVSS), an open industry standard for assessing the severity of computer system security vulnerabilities.
  • Threat Analysis and Risk Assessment (TARA) methodology from the ISO/SAE 21434:2021 ("Road Vehicles - Cybersecurity Engineering") Standard, jointly developed by the International Organisation for Standardisation (ISO) and the Society of Automotive Engineers (SAE), for managing cybersecurity threats and vulnerabilities throughout the life cycle of a vehicle.
To demonstrate the practical application of ACTISM, we developed a comprehensive security model for an IVI system. We outline the step-by-step process of building the security model, emphasising the integration of threat intelligence and contextual information to accurately characterise and assess cybersecurity risks. Furthermore, we illustrate ACTISM’s adaptive response to a real-world vulnerability disclosure affecting the IVI system, underscoring its effectiveness in addressing emerging threats and showcasing its potential applicability to other automotive systems.
Finally, we outline avenues for future research and development in this area.
Related Work
Various articles have attempted to address security modelling in automotive systems.
Das et al. [2] employed the STRIDE methodology for threat modelling and utilised the SAHARA (Safety-Aware Hazard Analysis and Risk Assessment) and DREAD (Damage, Reproducibility, Exploitability, Affected Users, and Discoverability) methodologies to assess risks associated with Infotainment High Performance Computing (HPC) systems. However, their analysis did not incorporate attack modelling.
Abuabed et al. [3] proposed a framework for assessing the cybersecurity vulnerabilities of automotive vehicles. Their approach employed STRIDE to classify threats. For assessing the impact of threats, they used version 1.0 of the HEAVENS security model [4], and they evaluated attack feasibility with CVSS version 2.0. Additionally, they used a risk matrix to assess risks based on a use-case involving the Advanced Driver Assistance System (ADAS). Using a TARA approach compatible with the ISO/SAE 21434:2021 standard, the framework utilises attack trees to determine attack paths, facilitating a focused analysis of relevant threat vectors and potential attacks during the assessment of attack feasibility.
The framework has certain limitations. Firstly, the threat identification process did not include physical attack vectors, such as accessing and modifying embedded software in vehicle components, in the STRIDE Data Flow Diagram (DFD). This may arise from the assumption that physical attacks do not involve data flows. However, physical attack vectors, such as "USB Man-in-the-Middle (USB MITM)" attacks, can exploit and facilitate data flows by intercepting and altering communication. As a result, neglecting physical attack vectors introduces a significant gap in the overall security design of the vehicle.
Secondly, both the HEAVENS version 1.0 and CVSS version 2.0 frameworks adopted in Abuabed et al. [3] are no longer current. Their newer versions include improvements that deliver more granularity and accuracy in the resulting risk assessment. Lastly, the framework does not adequately address threats in later stages of the life cycle. As automotive system life cycles are typically 10 to 20 years or longer, the vehicle system could become increasingly vulnerable to new types of attacks that were not anticipated during the initial security assessment.
Chlup et al. [5] proposed an automated threat modelling tool for the automotive domain. It combines threat modelling and attack tree analysis using automatically-generated attack trees and attack paths. Additionally, it uses anti-patterns to perform in-depth analysis in the threat model. However, the approach does not perform attack feasibility estimation and risk assessment, which are key steps in the TARA workflow.
Other literature relevant to automotive cybersecurity includes a study by Girdhar et al. [6], who used the STRIDE methodology to create a threat model focused on post-accident cyberattack event analysis for CAV.
Additionally, Dang et al. [7] used a STRIDE framework to analyse and classify cybersecurity threats in an autonomous vehicle architecture, including the infotainment system. While the authors ranked the threats using CVSS version 3.0 scores, they did not assess the corresponding impact. The authors focused only on cybersecurity threats originating from outside the vehicle. This means that attacks that involve access to the vehicle’s internal architecture from within the vehicle are excluded in the analysis.
The current landscape of security modelling highlights several limitations. Firstly, there are differing assumptions regarding which vehicle assets fall within the trust boundaries of security models. Secondly, the feasibility and impact of cyber-physical attacks are rarely considered in security modelling. These gaps underscore the need for an iterative and cohesive framework to effectively address the evolving challenges in automotive cybersecurity.
Threat-informed Dynamic Security Modelling with ACTISM
We propose ACTISM (Automotive Consequence-Driven and Threat-Informed Security Modelling) - a methodology for modelling security threats and attack vectors in a dynamic, iterative manner, specifically designed for automotive systems. We will illustrate the application of ACTISM to an IVI system as a running example.
Unlike conventional security modelling methodologies, ACTISM seamlessly integrates both threat and attack modelling in a continuous, iterative workflow throughout all stages of the vehicle’s life cycle. Moreover, ACTISM is dynamic, receiving updates through cybersecurity threat information acquired from vulnerability advisories and ongoing security monitoring.
Figure 1: Overview of threat-informed dynamic security modelling
As illustrated in Figure 1, a threat-informed dynamic security modeling approach relies on the seamless integration of threat and attack modeling, utilising methodologies and tools like Microsoft’s STRIDE and CVSS 3.1. A key component of this approach is the strategic incorporation of threat intelligence through continuous monitoring, which ensures that the security model remains adaptive and effective in mitigating emerging risks.
Such an iterative modelling approach is vital because threat and attack models created in the early stages of system life cycles can become outdated or irrelevant over time as new vulnerabilities emerge. This gap widens, particularly when cyber-physical attacks occur like those targeting keyless entry systems [8], breaching the trust boundary within security models.
ACTISM does not replace threat and attack modelling. Instead, it streamlines their adoption and utilisation by cybersecurity practitioners. By incorporating an iterative modelling approach, the methodology enhances the practicality and efficacy of addressing evolving cybersecurity threats and attacker capabilities. ACTISM enhances the TARA framework by dynamically updating the cybersecurity posture of automotive systems based on the evolving threat landscape and the specific impacts of potential attacks on system functionality and safety. ACTISM comprises eight key steps, outlined as follows:
  1. 1.
    Identify assets
  2. 2.
    Construct Data Flow Diagrams
  3. 3.
    Model threats
  4. 4.
    Evaluate impact
  5. 5.
    Model attacks
  6. 6.
    Evaluate attack feasibility
  7. 7.
    Evaluate risks
  8. 8.
    Review, update, and iterate
The details of each step will be elaborated on in the next section, using the running example of an IVI system.
Lastly, we further incorporate the following enhancements in ACTISM to increase its practical utility amongst cybersecurity practitioners:
  • Threat analysis: We employ the STRIDE methodology using a STRIDE-per-element approach where threats are assessed against each element in the Data Flow Diagram. We do not use a STRIDE-per-interaction approach, as this yields a statistically significant lower level of completeness compared to the STRIDE-per-element approach [9]. Additionally, we use HEAVENS version 2.0 [10] for impact assessment, as version 2.0 offers enhancements over version 1.0, providing more accurate and comprehensive risk assessments.
  • Attack analysis: We use CVSS version 3.1 to assess attack feasibility ratings, allowing us to leverage current cybersecurity vulnerability knowledge bases. Additionally, CVSS version 3.1 offers granular metrics and scoring options, which consider factors affecting the exploitability and impact of a vulnerability.
Applying ACTISM to an In-Vehicle Infotainment (IVI) System
To illustrate the practical application of ACTISM, it is essential to understand the IVI system and its role in automotive cybersecurity.
Legacy head units offered basic features such as radio reception and CD playback, with limited connectivity. In contrast, modern IVI systems are advanced digital platforms that integrate various functionalities beyond audio and navigation. These systems serve as central hubs, providing seamless connectivity to external networks, managing complex data streams, enabling access to vehicle diagnostics, and supporting interactive user interfaces for both drivers and passengers.
While modern IVI systems enhance convenience and operational efficiency, they also expose vehicles to increased cybersecurity risks. These systems are vulnerable to traditional threats, such as remote exploits and data breaches, as well as emerging threats like cyber-physical attacks targeting vehicle functionalities.
The complexity of modern IVI systems significantly expands their attack surface. They integrate multiple wireless communication protocols—Wi-Fi, Bluetooth, and cellular networks—and connect to external services and mobile devices, increasing potential vulnerabilities. Malicious actors can exploit weaknesses in wireless protocols or insecure network connections to gain unauthorised access. Furthermore, advanced functionalities like voice recognition and artificial intelligence add layers of complexity, creating additional attack vectors.
A dynamic approach like ACTISM allows for iterative security modelling throughout the IVI system’s life cycle, ensuring comprehensive protection against evolving threats.
Step 1: Identify assets
In the first step, identifying vehicle assets within the IVI system threat model typically involves a collaborative effort among system architects, engineers, and cybersecurity experts. This process integrates various activities, including reviewing technical specifications, examining system architecture diagrams, and analysing functional requirements documents. Additionally, considering the interconnected nature of modern IVI systems, potential dependencies and interactions between assets are also analysed to understand how vulnerabilities in one component could impact overall system security. These efforts aim to ensure a comprehensive understanding and coverage of all relevant assets within the IVI system.
In this article, we reference the IVI HPC system DFD from Das et al. (2024) [2] to guide our identification of IVI system assets. This initial asset identification, listed below, serves as a foundational step for conducting further detailed analysis and security modelling.
  • In-vehicle modules
    • In-Vehicle Infotainment (IVI) System
    • Human-machine Interface (HMI)
    • Central Gateway
    • GPS Receiver
    • USB
    • Bluetooth
    • Wi-Fi
  • Data stores
    • Local maps data store
    • IVI system embedded data store
Step 2: Construct Data Flow Diagrams
In the next step, we construct and analyse DFDs to identify the vehicle trust boundary and potential threat and attack vectors. This process uses the Microsoft Threat Modelling Tool (Version 7.3.31026.3) along with the Automotive Threat Modelling Template [11] developed by the NCC Group. We conduct DFD analysis for both a typical automotive system [12] and specifically for the IVI system (Figure 2) to ensure thorough identification and analysis of potential threats and vulnerabilities.
Figure 2: Data Flow Diagram for an In-Vehicle Infotainment (IVI) System
Step 3: Model threats
Next, we use the Microsoft Threat Modelling Tool to generate threats based on the previously constructed Data Flow Diagram (DFD) for the IVI system. The tool categorises these threats according to the STRIDE threat model automatically, providing an initial classification into categories such as Spoofing, Tampering, Repudiation, Information Disclosure, Denial of Service, and Elevation of Privilege.
While the tool categorises threats by default and attempts to provide threat descriptions based on its own threat catalogues and templates, the initial output may be somewhat generic and not specifically tailored to the nuances of the IVI system’s operating environment. Therefore, our analysis involves further contextualisation and systematic review of these categorised threats.
We focus on the in-vehicle modules and data stores identified in the earlier step, ensuring that the generated threats are relevant to these components. After refining the list, we update threat descriptions to better align with the specific interfaces and nuances of the IVI system. A domain expert can perform this refinement relatively easily using the Microsoft Threat Modelling Tool, taking about 30 minutes for this example. The final curated list of threats for the IVI system is shown in Figure 3.
Figure 3: Threats generated by Microsoft Threat Modelling Tool
Step 4: Evaluate impact
To analyse the impact of the identified threats, we use the HEAVENS version 2.0 framework [10] which is specifically designed for automotive risk assessments. This framework allows us to systematically evaluate the potential impacts across four key dimensions: safety (s), financial (f), operational (o), and privacy and legislation (pl).
Our approach to impact analysis is as follows:
  • Assess impact values: We assign numerical values to each threat based on its impact level on safety, financial stability, operational continuity, and compliance with privacy and legislation. (Table 1)
    Table 1: Mapping of impact levels to numerical values
    Impact Level Value
    None 0
    Low 1
    Medium 10
    High 100
  • Calculate impact score: Equation 1 is used to compute the overall impact score of each threat. This equation combines the weighted values of safety, financial, operational, and privacy and legislation impacts. To reflect the critical importance of safety in the context of automotive cybersecurity, we set $w_{s}=10$, placing the maximum weight on safety.
    Impact $\small\text{Impact Score}=\frac{w_{s}i_{s}+w_{f}i_{f}+w_{o}i_{o}+w_{pl}i_{pl}}{w_{s}+w_{f}+w_{o}+w_{pl}}$ (1)
    where $w_{s}=10$ and $w_{f}=w_{o}=w_{pl}=1$.
  • Derive impact rating: Based on the computed impact score of each threat, we derive the corresponding ratings using the thresholds defined in Table 2. This rating helps prioritise threats based on their potential severity and guides subsequent risk mitigation in our analysis.
    Table 2: Mapping of computed impact scores to ratings
    Impact Score Rating
    0.00 None
    $0.010.00<x<$ 0.01 Low
    $0.050.01\leq x<$ 0.05 Medium
    $0.450.05\leq x<$ 0.45 High
    $1.000.45\leq x<$ 1.00 Critical
The final impact evaluation is tabulated in Figure 4.
Figure 4: Impact assessment based on HEAVENS 2.0
Step 5: Model attacks
In this step, we use an attack tree to model attacker goals and map out attack paths associated with threat and attack vectors identified earlier using the Microsoft Threat Modelling Tool.
Since the vehicle driver is within the car trust boundary (Figure 2), we will assume the attacker personas and their goals as follows:
  • Persona A: A car thief interested in stealing sensitive data from the In-Vehicle Infotainment system.
  • Persona B: A rogue security researcher interested in manipulating vehicle functions that compromise the safety of the vehicle.
The attack tree depicted in Figure 5 illustrates the potential attack paths these two personas could employ to achieve their respective goals.
Figure 5: Attack tree for IVI system
Step 6: Evaluate attack feasibility
Next, we use CVSS version 3.1 to analyse and compute the attack feasibility ratings. CVSS provides a standardised method for assessing the severity of vulnerabilities based on their impact and exploitability factors.
The CVSS 3.1 base score is computed using Equation 2.
Base Score=Round up(0.6×Impact+0.4×Exploitability1.5Impact $\tiny\text{Base Score}=\text{Round up}\left(\frac{{0.6\times\text{Impact}+0.4\times\text{Exploitability}-1.5}}{{\text{Impact Factor}+0.8}}\right)$ (2)
where Impact, Exploitability, and Impact Factor are derived from the following components:
  • Impact Metrics (C, I, A):
    • Confidentiality (C): The potential impact on confidentiality.
    • Integrity (I): The potential impact on integrity.
    • Availability (A): The potential impact on availability.
  • Exploitability Metrics (AV, AC, PR, UI):
    • Attack Vector (AV): How the vulnerability is accessed.
    • Attack Complexity (AC): The complexity of the attack required to exploit the vulnerability.
    • Privileges Required (PR): The level of privileges required to exploit the vulnerability.
    • User Interaction (UI): The level of user interaction required to exploit the vulnerability.
  • Impact Factor:
    • The Impact Factor depends on the scope of the vulnerability, which can be either Scope Unchanged (U) or Scope Changed (S). If the vulnerability affects other system components, the Scope is considered Changed (S), otherwise, it is Unchanged (U).
Additionally, the CVSS temporal score takes into account factors such as exploit code maturity, remediation level, and report confidence. It is calculated using Equation 3.
Temporal Score=Round up(Base Score×Exploit Code Maturity×Remediation Level×Report $\tiny\begin{split}\text{Temporal Score}=\text{Round up}(\text{Base Score}\times\text{Exploit Code Maturity}\times\\ \text{Remediation Level}\times\text{Report Confidence})\end{split}$ (3)
Figure 6: Attack feasibility ratings based on CVSS version 3.1
Figure 6 lists the attack feasibility ratings computed using the CVSS version 3.1 calculator (https://nvd.nist.gov/vuln-metrics/cvss/v3-calculator). These ratings are based on the information available at the time of assessment during the early design stage of the IVI system.
Figure 7: Risk ratings evaluated using a 5x5 risk matrix
Step 7: Evaluate risks
In this step, we evaluate the cybersecurity risks of the IVI system using a 5x5 risk matrix. The corresponding risk ratings for identified threats are tabulated in Figure 7.
These risk ratings serve as a guideline for the Original Equipment Manufacturer (OEM) of the vehicle or the manufacturer of the IVI system. They can use these ratings to prioritise the implementation of mitigation controls during the design phase of the system life cycle. This approach ensures that adequate attention and resources are allocated to address the most critical cybersecurity concerns identified through this comprehensive analysis.
Step 8: Review, update, and iterate
In this final step, we continuously review the cybersecurity analysis and update the risk assessment based on new threat information, insights or changes in the automotive system environment. The updated risk assessment helps refine existing mitigation strategies and enhances security controls against emerging cybersecurity threats. This iterative cycle of review and improvement is crucial for maintaining the automotive system’s long-term cybersecurity posture.
Updating the Security Model After a Vulnerability Disclosure
In the earlier section, we discussed the TARA framework as part of our approach to security modelling for automotive systems. Generally, this is performed during the early stages of system life cycles. However, evolving cybersecurity threats necessitate the incorporation of new knowledge gained from ongoing security monitoring and threat intelligence throughout the system’s life cycle.
ACTISM can be seamlessly integrated with the TARA framework to create a threat-informed dynamic security modelling methodology.
We now apply this methodology to a real-world example using the Tesla Electric Vehicle (EV) jail-break vulnerability disclosure [13].
After gathering technical details from the vulnerability disclosure, we review the IVI system security model developed in the earlier section and update the CVSS metrics as necessary.
For this vulnerability disclosure, there is no need to include additional assets in the security model or to revise the DFD (Figure 2) and attack tree (Figure 5). Although there is an opportunity to enhance the DFD and attack tree to include new attack vectors, such as cyber-physical attacks like keyless entry systems [8], we will not do so in this example. However, it is critical to reassess the CVSS metrics associated with Threat ID 9146, which correlates with the vulnerability disclosure.
The CVSS metrics corresponding to Threat ID 9146 are updated as follows:
  • Scope (S) From Unchanged (UC) to Changed (C), as successful exploitation of the vulnerability could grant an attacker full control over the IVI system, potentially enabling them to access and manipulate certain vehicle functions.
  • Integrity (I) From Low (L) to High (H), as successful exploitation of the vulnerability could grant an attacker root-level read and write access to the IVI file system.
  • Availability (A) From None (N) to High (H), as an attacker could deny access to resources in the IVI system.
  • Exploit Code Maturity (E) From Unproven (U) to Proof-of-Concept (P), in view of the disclosure from the security researchers.
  • Remediation Level (RL) From Unavailable (U) to Temporary Fix (T), as the AMD Trusted Platform Module (TPM) in affected IVI systems could still be susceptible to voltage glitching attacks.
  • Report Confidence (RC) From Unknown (U) to Confirmed (C), in view of the disclosure from the security researchers.
The updated risk rating based on these revised metrics is shown in Figure 8. In summary, the corresponding risk changed from Medium to High.
This ability to perform timely updates to the risk assessment is crucial for maintaining robust cybersecurity in automotive systems. By continuously adapting to new threats and changes in attack feasibility and impact, proactive measures can be implemented to safeguard automotive system safety and functionality. This dynamic approach not only enhances the resilience of automotive systems against evolving cybersecurity threats but also reinforces the importance of informed decision-making in the face of emerging risks. Ultimately, it empowers stakeholders to take appropriate actions that protect both the automotive system and its users.
Figure 8: Updated risk rating following vulnerability disclosure
Discussion
Unlike early-stage security modelling, where OEMs and IVI system manufacturers implement comprehensive measures, addressing cyber-physical vulnerabilities later in the life cycle often involves temporary fixes or workarounds. These may not fully mitigate the risks, so vehicle owners can use risk ratings to make informed decisions. By adopting additional precautions, they can enhance the cybersecurity resilience of their vehicles throughout their life cycle.
This underscores the need for a continuous, iterative, and threat-informed security modelling process. Ongoing threat intelligence gathering, including real-time monitoring and tracking vulnerability disclosures, is crucial. A comprehensive repository of vulnerability attributes, stored in a standardised format such as JSON, enables automation. Tools can parse this data, extract relevant information, and enrich the security model via workflows like TARA.
CVSS scores and risk ratings could then be dynamically recalculated based on updated threat context. Security models can be formalised using Domain-Specific Languages (DSLs) such as the Meta Attack Language (MAL), allowing clearer descriptions of threats and vulnerabilities. Furthermore, adversary tactics from the MITRE ATT&CK knowledge base could be incorporated to enhance these models.
Machine learning approaches could also help prioritise vulnerabilities by analysing their impact, using historical data and threat intelligence to improve automotive cybersecurity.
Conclusion
Automotive systems are increasingly facing cybersecurity threats, with attackers targeting vulnerabilities that pose risks to data, functionality, and safety. Ensuring the security of automotive systems is critical to maintaining their integrity and reliability.
We introduced ACTISM (Automotive Consequence-Driven and Threat-Informed Security Modelling), a novel approach designed to enhance the resilience of automotive systems by dynamically updating their cybersecurity posture. ACTISM responds to evolving threats and attacker tactics, ensuring security models remain relevant throughout a system’s lifecycle.
Unlike traditional models, ACTISM continuously integrates new data from threat intelligence and security monitoring, keeping protection up to date. It also captures the multi-agent nature of cyber-physical attacks, making it applicable to automotive domains. This integration strengthens the accuracy and completeness of security models, reflecting the complexity of real-world cyber-physical attacks.
Cybersecurity incidents in automotive systems differ from traditional IT breaches due to their potential physical consequences, such as vehicle misnavigation or compromised safety. ACTISM’s consequence-driven approach ensures that cyber-physical risks are considered, directly addressing the operational and safety integrity of automotive systems.
To demonstrate ACTISM’s practical application, we developed a comprehensive security model for an IVI system. This involved identifying key assets, conducting threat modelling, and performing risk assessments, resulting in a robust security framework. A real-world case study further highlighted ACTISM’s ability to adapt to emerging threats and its potential for broader use in other automotive technologies.
In conclusion, ACTISM offers a proactive, adaptable approach to automotive cybersecurity, addressing current challenges while preparing for the evolving landscape of automotive technology.

### v2

@@
-Article Type: Pre-Print
+ACTISM: Threat-Informed Dynamic Security Modelling for Automotive Systems
 Abstract
-Cybersecurity threats in automotive systems pose significant risks to safety and reliability.
-This article introduces a methodology integrating threat-informed dynamic security modelling with a Threat Analysis and Risk Assessment workflow.
-Using the example of an In-Vehicle Infotainment system, we demonstrate the methodology’s application in risk management to strengthen automotive resiliency.
-Keywords: risk management, security, autonomous vehicles
-journal: IEEE Security & Privacy
-Introduction
-Automotive systems have transformed into sophisticated platforms brimming with cutting-edge features.
+Cybersecurity threats in complex cyber-physical systems pose significant risks to system functionality and safety.
+This experience report introduces ACTISM (Automotive Consequence-Driven and Threat-Informed Security Modelling) — an integrated security modeling approach that enhances the resilience of automotive systems by dynamically updating their cybersecurity posture in response to prevailing threats, attacker tactics, and their impact on system functionality and safety.
+ACTISM addresses the existing knowledge gap in static security assessment methodologies by providing a dynamic and iterative framework.
+We demonstrate the effectiveness of ACTISM by applying it to a real-world example of the Tesla Electric Vehicle’s In-Vehicle Infotainment system, illustrating how the security model can be adapted as new threats emerge.
+We also outline avenues for future research and development in this area, including automated vulnerability management workflows for automotive systems.
+Keywords:
+Risk management Security Automotive systems
+1 Introduction
+Automotive systems have transformed into complex cyber-physical systems brimming with cutting-edge features.
 In contrast to their predecessors, today’s automotive systems are equipped with a variety of advanced sensors and connectivity options, such as Control Area Networks (CAN), Local Interconnect Networks (LIN), FlexRay, and Ethernet.
-They also offer seamless external connectivity through Bluetooth, Wi-Fi, and GPS.
-Additionally, the integration of feature-rich In-Vehicle Infotainment (IVI) systems and powerful computing units for autonomous driving continues to push the boundaries of what is possible in automotive technology.
+They also offer external connectivity through Bluetooth, Wi-Fi, and GPS.
+In addition, the integration of feature-rich In-Vehicle Infotainment (IVI) systems and powerful computing units for autonomous driving continues to push the boundaries of what is possible in automotive technology.
 However, this technological evolution also broadens the attack surface, making automotive systems more susceptible to cybersecurity attacks.
-Cybersecurity threats and vulnerabilities in the automotive industry are well-documented, and attackers exploit these vulnerabilities to compromise vehicle security, access sensitive data, and even take control of vehicle functions.
+Cybersecurity threats and vulnerabilities in the automotive industry are well documented, and attackers exploit these vulnerabilities to compromise vehicle security, access sensitive data, and even take control of vehicle functions.
 Recent literature has increasingly highlighted the challenges of defending against cybersecurity threats in automotive systems.
-In their survey on security attacks and defence techniques for Connected and Autonomous Vehicles (CAVs), Pham and Xiong (2021) observed that attackers could potentially compromise one vehicle component after another.
+In their survey on security attacks and defence techniques for Connected and Autonomous Vehicles (CAVs), Pham and Xiong [11] observed that attackers could potentially compromise one vehicle component after another.
 To mitigate this risk, they emphasised the need for manufacturers to strengthen the security of all system components to ensure the overall protection of CAVs.
@@
 Additionally, they enable consumers to make informed decisions when selecting vehicles based on their cybersecurity robustness.
-However, many of the threat and attack models discussed in the literature do not fully address the unique cybersecurity requirements of modern automotive systems.
+However, many of the threat and attack models discussed in the literature do not fully address the unique cybersecurity requirements of modern automotive systems. These requirements include:
+  • Interconnectivity: Modern vehicles are highly interconnected systems, comprising numerous Electronic Control Units (ECUs) that communicate with each other.
+Ensuring secure communication between these components is essential to prevent unauthorised access and control.
+  • Safety-Critical Systems: The integration of advanced driver-assistance systems (ADAS) and autonomous driving technologies means that any cybersecurity breach could directly impact vehicle safety.
+Ensuring the security of these systems is paramount to protecting the safety of passengers and other road users.
+  • Physical Security: Automotive systems are subject to physical threats.
+Attackers can potentially gain physical access to vehicles, enabling them to tamper with hardware components or inject malicious code directly.
+  • Over-the-Air (OTA) Updates: Many modern vehicles support OTA updates for software and firmware. While this feature enhances convenience and functionality, it also introduces new attack vectors.
+  • User Privacy: Automotive systems collect and process a significant amount of data, including personal information, driving habits, and location data.
+Protecting user privacy is crucial to prevent misuse of this data.
 In our previous work [7], we conducted a systematic literature review of threat and attack modelling approaches pertinent to cyber-physical systems (CPSs)—systems that integrate computational and physical process—of which automotive systems are an example.
@@
   • Ambiguity persists in the literature regarding the definitions and relationships between threat modelling and attack modelling.
+For instance, threat models developed using Microsoft’s STRIDE (Spoofing, Tampering, Repudiation, Information Disclosure, Denial of Service, and Elevation of Privilege) primarily focus on data flows and information security.
+However, STRIDE was not designed with automotive systems in mind, making its application in this context somewhat limited.
+Consequently, STRIDE often overlooks critical aspects of how cybersecurity breaches can directly affect automotive functions and safety.
+As a result, the reliability of the general outcomes from STRIDE is questionable when applied to the unique characteristics of automotive environments.
+In other words, threat modelling of automotive systems often does not consider the consequences of cybersecurity attacks beyond information security.
+Similarly, attack models designed for traditional IT systems do not fully align with the multi-path tactics and techniques employed by attackers seeking to gain control of a vehicle.
+These techniques often involve a combination of physical and cyber actions, necessitating a more comprehensive and dynamic approach to security modelling.
 These findings highlight the need to tailor security modeling approaches for automotive systems so that they can adequately capture the unique cybersecurity threats faced in these environments.
-For instance, threat models developed using Microsoft’s STRIDE (Spoofing, Tampering, Repudiation, Information Disclosure, Denial of Service, and Elevation of Privilege) primarily focus on data flows and information security, overlooking critical aspects of how such breaches can directly affect automotive functions and safety.
-In other words, threat modelling of automotive systems often do not consider consequences of cybersecurity attacks beyond information security.
-On a similar note, attack models designed with traditional IT systems in mind do not fully align with multi-path tactics and techniques employed by attackers seeking to gain control of a vehicle.
-These techniques often involve a combination of physical and cyber actions.
-In short, our investigation has highlighted limitations in current threat and attack modelling methods that underscore the need for a tailored, dynamic approach.
-Threat and attack models need to iteratively and continuously consider both consequences and characteristics of cyber-physical attacks so as to offer a comprehensive approach to safeguarding automotive systems.
 Contributions.
-In this article, we introduce ACTISM (Automotive Consequence-Driven and Threat-Informed Security Modelling) — a novel security modeling approach that enhances the resilience of automotive systems by dynamically updating their cybersecurity posture in response to prevailing threats, attacker tactics, and their impact on system functionality and safety.
+In this experience report, we introduce ACTISM (Automotive Consequence-Driven and Threat-Informed Security Modelling) — an integrated security modeling approach that enhances the resilience of automotive systems by dynamically updating their cybersecurity posture in response to prevailing threats, attacker tactics, and their impact on system functionality and safety.
+This approach addresses the existing knowledge gap in static security assessment methodologies by providing a dynamic and iterative framework.
 We demonstrate the effectiveness of ACTISM by applying it to an IVI system, illustrating how the model can be adapted as new threats emerge.
-Unlike existing security modelling approaches, ACTISM dynamically updates security models with new information acquired through threat intelligence and ongoing security monitoring.
+Unlike existing security modelling approaches, ACTISM integrates various frameworks to dynamically update security models with new information, acquired through threat intelligence and ongoing security monitoring.
 This keeps security models relevant against evolving attacker tactics, techniques and procedures (TTPs) even if the models were developed in early stages of system development.
 ACTISM also allows practitioners to apply the security models in both IT and automotive domains by using attack paths to capture the multi-path and multi-agent aspects of real-world cyber-physical attacks.
-Cybersecurity incidents in automotive systems differ significantly from those in traditional IT systems.
-These incidents can lead to complex failure modes with both cyber and physical consequences, such as vehicle misnavigation and potentially dangerous outcomes.
-By adopting a consequence-driven and cyber-informed approach to cybersecurity, ACTISM ensures that cyber-physical attacks, effects and consequences are considered in security modelling for automotive systems.
-ACTISM seamlessly and systematically integrates with, and adapts to a number of threat analysis and risk assessment frameworks such as:
+In order to bolster the completeness of the security model, ACTISM integrates the following threat analysis and risk assessment frameworks:
   • Microsoft’s STRIDE methodology.
-  • HEAVENS (HEAling Vulnerabilities to ENhance Software Security and Safety), a structured and model-based hazard analysis and risk assessment method for automotive systems.
+  • HEAVENS (HEAling Vulnerabilities to ENhance Software Security and Safety), a structured and model-based hazard analysis and risk assessment method specifically designed for automotive systems.
   • Attack Tree, a hierarchical diagram representing potential attack paths, illustrating how an attacker could exploit system vulnerabilities to achieve specific goals.
   • Common Vulnerability Scoring System (CVSS), an open industry standard for assessing the severity of computer system security vulnerabilities.
+CVSS (Common Vulnerability Scoring System): This model is also static as it provides a standardised method for evaluating the severity of vulnerabilities at a particular point in time.
+It does not dynamically adapt to new threats without manual intervention.
   • Threat Analysis and Risk Assessment (TARA) methodology from the ISO/SAE 21434:2021 ("Road Vehicles - Cybersecurity Engineering") Standard, jointly developed by the International Organisation for Standardisation (ISO) and the Society of Automotive Engineers (SAE), for managing cybersecurity threats and vulnerabilities throughout the life cycle of a vehicle.
+Table 1 lists the differences between ACTISM and these security modelling frameworks, and their respective limitations.
+In contrast with other security modelling frameworks, the ACTISM framework integrates various frameworks to be dynamic, continuously adapting to new threats and updating the automotive system’s cybersecurity posture.
+This distinction sets it apart from the static models, providing a more robust approach to maintaining automotive cybersecurity.
+Table 1: Comparison of ACTISM with other security modelling frameworks
+Framework Type Limitations
+ACTISM Dynamic Relies on the integration of various security modelling frameworks.
+STRIDE Static Limited to specific threat categories. Does not dynamically adapt to new threats without manual intervention.
+Attack Trees Static Requires manual updates and may not cover all emerging threats. Does not dynamically adapt to new threats without manual intervention.
+HEAVENS Static Primarily focuses on high-level security analysis which may not cover detailed, system-specific vulnerabilities and threats.
+Does not dynamically adapt to new threats without manual intervention.
+CVSS Static Does not dynamically adapt to new threats without manual intervention.
+TARA Static Does not dynamically adapt to new threats without manual intervention.
+Although frameworks such as STRIDE and CVSS were designed for IT security threats, they can also be applied to physical threats.
+As an example, tampering with physical components can lead to unauthorised access, data manipulation, or service disruption.
+As an integrative approach, ACTISM provides guidance on how both cyber and physical attacks can be modeled using these frameworks to capture the various entry points and methods attackers might use, ensuring comprehensive security coverage across both cyber and physical domains.
 To demonstrate the practical application of ACTISM, we developed a comprehensive security model for an IVI system.
 We outline the step-by-step process of building the security model, emphasising the integration of threat intelligence and contextual information to accurately characterise and assess cybersecurity risks.
-Furthermore, we illustrate ACTISM’s adaptive response to a real-world vulnerability disclosure affecting the IVI system, underscoring its effectiveness in addressing emerging threats and showcasing its potential applicability to other automotive systems.
-Finally, we outline avenues for future research and development in this area.
-Related Work
-Various articles have attempted to address security modelling in automotive systems.
+Furthermore, we demonstrate the utility of ACTISM by illustrating its adaptive response to a real-world vulnerability disclosure affecting the IVI system.
+This case study underscored ACTISM’s effectiveness in addressing emerging threats and evolving security challenges in automotive cybersecurity.
+Furthermore, the study showcased the versatility of ACTISM, demonstrating its potential applicability to other automotive systems beyond the IVI, making it a robust tool for safeguarding a wide range of vehicle technologies.
+Finally, we outline avenues for future research and development in this area, including automated vulnerability management workflows.
+2 Related Work
+Various works have attempted to address security modelling in automotive systems.
 Das et al.
@@
 Their approach employed STRIDE to classify threats.
-For assessing the impact of threats, they used version 1.0 of the HEAVENS security model [4], and they evaluated attack feasibility with CVSS version 2.0.
+For assessing the impact of threats, they used version 2.0 of the HEAVENS security model [9], and they evaluated attack feasibility with CVSS version 2.0.
 Additionally, they used a risk matrix to assess risks based on a use-case involving the Advanced Driver Assistance System (ADAS).
-Using a TARA approach compatible with the ISO/SAE 21434:2021 standard, the framework utilises attack trees to determine attack paths, facilitating a focused analysis of relevant threat vectors and potential attacks during the assessment of attack feasibility.
-The framework has certain limitations.
-Firstly, the threat identification process did not include physical attack vectors, such as accessing and modifying embedded software in vehicle components, in the STRIDE Data Flow Diagram (DFD).
-This may arise from the assumption that physical attacks do not involve data flows.
-However, physical attack vectors, such as "USB Man-in-the-Middle (USB MITM)" attacks, can exploit and facilitate data flows by intercepting and altering communication.
-As a result, neglecting physical attack vectors introduces a significant gap in the overall security design of the vehicle.
-Secondly, both the HEAVENS version 1.0 and CVSS version 2.0 frameworks adopted in Abuabed et al.
-[3] are no longer current.
-Their newer versions include improvements that deliver more granularity and accuracy in the resulting risk assessment.
+Using a TARA approach compatible with the ISO/SAE 21434:2021 standard, the framework utilises attack trees to determine attack paths.
+This method facilitates a focused analysis of relevant threat vectors and potential attacks during the assessment of attack feasibility.
+However, the framework has several limitations.
+Firstly, in the threat identification process, the authors did not include physical attack vectors, such as accessing and modifying embedded software in vehicle components, in the STRIDE Data Flow Diagram (DFD).
+This omission could be because of the assumption that physical attack vectors do not involve flows or exchanges of data.
+Notwithstanding this assumption, physical attack vectors can be exploited to facilitate data flows, such as "USB Man-in-the-Middle (USB MITM)" attacks, where an attacker intercepts and potentially alters the communication between a USB device and a host system.
+Consequently, physical attack vectors can be a crucial gap in the overall security design of the vehicle.
+Secondly, the CVSS version 2.0 framework adopted in Abuabed et al.
+[1] is no longer current.
+The newer versions of CVSS include improvements that deliver more granularity and accuracy in the resulting risk assessment.
 Lastly, the framework does not adequately address threats in later stages of the life cycle.
@@
 It combines threat modelling and attack tree analysis using automatically-generated attack trees and attack paths.
-Additionally, it uses anti-patterns to perform in-depth analysis in the threat model.
+In addition, it uses anti-patterns to perform in-depth analysis in the threat model.
 However, the approach does not perform attack feasibility estimation and risk assessment, which are key steps in the TARA workflow.
 Other literature relevant to automotive cybersecurity includes a study by Girdhar et al.
-[6], who used the STRIDE methodology to create a threat model focused on post-accident cyberattack event analysis for CAV.
+[5] who used the STRIDE methodology to create a threat model focused on post-accident cyberattack event analysis for CAV.
 Additionally, Dang et al.
@@
 These gaps underscore the need for an iterative and cohesive framework to effectively address the evolving challenges in automotive cybersecurity.
-Threat-informed Dynamic Security Modelling with ACTISM
-We propose ACTISM (Automotive Consequence-Driven and Threat-Informed Security Modelling) - a methodology for modelling security threats and attack vectors in a dynamic, iterative manner, specifically designed for automotive systems.
-We will illustrate the application of ACTISM to an IVI system as a running example.
-Unlike conventional security modelling methodologies, ACTISM seamlessly integrates both threat and attack modelling in a continuous, iterative workflow throughout all stages of the vehicle’s life cycle.
-Moreover, ACTISM is dynamic, receiving updates through cybersecurity threat information acquired from vulnerability advisories and ongoing security monitoring.
-Figure 1: Overview of threat-informed dynamic security modelling
-As illustrated in Figure 1, a threat-informed dynamic security modeling approach relies on the seamless integration of threat and attack modeling, utilising methodologies and tools like Microsoft’s STRIDE and CVSS 3.1.
-A key component of this approach is the strategic incorporation of threat intelligence through continuous monitoring, which ensures that the security model remains adaptive and effective in mitigating emerging risks.
-Such an iterative modelling approach is vital because threat and attack models created in the early stages of system life cycles can become outdated or irrelevant over time as new vulnerabilities emerge.
+3 Threat-informed Dynamic Security Modelling with ACTISM
+We propose ACTISM - a methodology for modelling security threats and attack vectors in a dynamic, iterative manner, specifically designed for automotive systems.
+We will illustrate the application of ACTISM to an IVI system as a running example and apply ACTISM to a real-world case study, where we demonstrate its ability to dynamically adapt to new threats and improve the security of automotive systems.
+This case study provides concrete evidence of the framework’s applicability and benefits, distinguishing our research from previous work.
+Current security modelling frameworks offer valuable insights but lack adaptability to evolving cyber-physical threats.
+ACTISM addresses this gap by introducing a framework that continuously updates the cybersecurity posture of automotive systems in response to both cyber and physical threats, attack vectors, and system vulnerabilities.
+Unlike static models, ACTISM incorporates threat intelligence through continuous monitoring, which ensures that the security model remains adaptive and effective in mitigating emerging risks.
+This iterative approach is vital because threat and attack models created in the early stages of system life cycles can become outdated or irrelevant over time as new vulnerabilities emerge.
 This gap widens, particularly when cyber-physical attacks occur like those targeting keyless entry systems [15], breaching the trust boundary within security models.
-ACTISM does not replace threat and attack modelling.
+ACTISM does not replace current threat and attack modelling frameworks.
 Instead, it streamlines their adoption and utilisation by cybersecurity practitioners.
-By incorporating an iterative modelling approach, the methodology enhances the practicality and efficacy of addressing evolving cybersecurity threats and attacker capabilities.
-ACTISM enhances the TARA framework by dynamically updating the cybersecurity posture of automotive systems based on the evolving threat landscape and the specific impacts of potential attacks on system functionality and safety.
-ACTISM comprises eight key steps, outlined as follows:
-  1. 1.
-    Identify assets
-  2. 2.
-    Construct Data Flow Diagrams
-  3. 3.
-    Model threats
-  4. 4.
-    Evaluate impact
-  5. 5.
-    Model attacks
-  6. 6.
-    Evaluate attack feasibility
-  7. 7.
-    Evaluate risks
-  8. 8.
-    Review, update, and iterate
-The details of each step will be elaborated on in the next section, using the running example of an IVI system.
-Lastly, we further incorporate the following enhancements in ACTISM to increase its practical utility amongst cybersecurity practitioners:
-  • Threat analysis: We employ the STRIDE methodology using a STRIDE-per-element approach where threats are assessed against each element in the Data Flow Diagram.
-We do not use a STRIDE-per-interaction approach, as this yields a statistically significant lower level of completeness compared to the STRIDE-per-element approach [9].
-Additionally, we use HEAVENS version 2.0 [10] for impact assessment, as version 2.0 offers enhancements over version 1.0, providing more accurate and comprehensive risk assessments.
-  • Attack analysis: We use CVSS version 3.1 to assess attack feasibility ratings, allowing us to leverage current cybersecurity vulnerability knowledge bases.
-Additionally, CVSS version 3.1 offers granular metrics and scoring options, which consider factors affecting the exploitability and impact of a vulnerability.
-Applying ACTISM to an In-Vehicle Infotainment (IVI) System
+By integrating various security modelling frameworks into an iterative approach, ACTISM enhances the practicality and efficacy of addressing evolving cybersecurity threats and attacker capabilities.
+ACTISM comprises eight key steps as shown in Figure 1.
+Figure 1: Overview of ACTISM
+To enhance the practical utility of ACTISM among cybersecurity practitioners, we adopt the STRIDE methodology using a STRIDE-per-element approach, where threats are assessed against each element in the Data Flow Diagram.
+We choose not to use the STRIDE-per-interaction approach, as research has shown it yields a statistically significant lower level of completeness compared to the STRIDE-per-element approach [13].
+Furthermore, we utilise HEAVENS version 2.0 [9] for impact assessment.
+To our knowledge, HEAVENS is the only framework designed to be fully compliant with the latest automotive cybersecurity standards, such as ISO/SAE 21434 and UNECE Regulation 155.
+In addition, we employ CVSS version 3.1 to assess attack feasibility, which allows us to leverage contemporary cybersecurity vulnerability knowledge bases.
+CVSS version 3.1 provides granular metrics and scoring options, considering various factors that affect the exploitability and impact of a vulnerability.
+This comprehensive approach ensures that ACTISM remains robust and adaptive to the dynamic nature of cybersecurity threats in automotive systems.
+4 Applying ACTISM to an In-Vehicle Infotainment (IVI) System
 To illustrate the practical application of ACTISM, it is essential to understand the IVI system and its role in automotive cybersecurity.
@@
 In contrast, modern IVI systems are advanced digital platforms that integrate various functionalities beyond audio and navigation.
-These systems serve as central hubs, providing seamless connectivity to external networks, managing complex data streams, enabling access to vehicle diagnostics, and supporting interactive user interfaces for both drivers and passengers.
+These systems serve as central hubs, providing connectivity to external networks, managing complex data streams, enabling access to vehicle diagnostics, and supporting interactive user interfaces for both drivers and passengers.
 While modern IVI systems enhance convenience and operational efficiency, they also expose vehicles to increased cybersecurity risks.
@@
 A dynamic approach like ACTISM allows for iterative security modelling throughout the IVI system’s life cycle, ensuring comprehensive protection against evolving threats.
-Step 1: Identify assets
+4.1 Step 1: Identify assets
 In the first step, identifying vehicle assets within the IVI system threat model typically involves a collaborative effort among system architects, engineers, and cybersecurity experts.
@@
 These efforts aim to ensure a comprehensive understanding and coverage of all relevant assets within the IVI system.
-In this article, we reference the IVI HPC system DFD from Das et al.
+In this experience report, we reference the IVI HPC system DFD from Das et al.
 (2024) [4] to guide our identification of IVI system assets.
 This initial asset identification, listed below, serves as a foundational step for conducting further detailed analysis and security modelling.
+It is important to note that this DFD is an example and may not encompass newer types of attacks, such as dirty road attacks that exploit the fact that autonomous vehicles rely on sensors and cameras to interpret their environment.
+Attackers may place objects or patterns on the road that confuse these sensors, leading to incorrect navigation decisions or even accidents.
   • In-vehicle modules
@@
     • IVI system embedded data store
-Step 2: Construct Data Flow Diagrams
+4.2 Step 2: Construct Data Flow Diagrams
 In the next step, we construct and analyse DFDs to identify the vehicle trust boundary and potential threat and attack vectors.
@@
 Figure 2: Data Flow Diagram for an In-Vehicle Infotainment (IVI) System
-Step 3: Model threats
+4.3 Step 3: Model threats
 Next, we use the Microsoft Threat Modelling Tool to generate threats based on the previously constructed Data Flow Diagram (DFD) for the IVI system.
@@
 A domain expert can perform this refinement relatively easily using the Microsoft Threat Modelling Tool, taking about 30 minutes for this example.
-The final curated list of threats for the IVI system is shown in Figure 3.
-Figure 3: Threats generated by Microsoft Threat Modelling Tool
-Step 4: Evaluate impact
-To analyse the impact of the identified threats, we use the HEAVENS version 2.0 framework [10] which is specifically designed for automotive risk assessments.
-This framework allows us to systematically evaluate the potential impacts across four key dimensions: safety (s), financial (f), operational (o), and privacy and legislation (pl).
+The final curated list of threats for the IVI system is shown in Table 9 [8].
+4.4 Step 4: Evaluate impact
+To evaluate the impact of the identified threats, we use the HEAVENS version 2.0 framework [9] which is specifically designed for automotive risk assessments.
 Our approach to impact analysis is as follows:
-  • Assess impact values: We assign numerical values to each threat based on its impact level on safety, financial stability, operational continuity, and compliance with privacy and legislation.
-(Table 1)
-    Table 1: Mapping of impact levels to numerical values
-    Impact Level Value
-    None 0
-    Low 1
-    Medium 10
-    High 100
-  • Calculate impact score: Equation 1 is used to compute the overall impact score of each threat.
-This equation combines the weighted values of safety, financial, operational, and privacy and legislation impacts.
-To reflect the critical importance of safety in the context of automotive cybersecurity, we set $w_{s}=10$, placing the maximum weight on safety.
-    Impact $\small\text{Impact Score}=\frac{w_{s}i_{s}+w_{f}i_{f}+w_{o}i_{o}+w_{pl}i_{pl}}{w_{s}+w_{f}+w_{o}+w_{pl}}$ (1)
-    where $w_{s}=10$ and $w_{f}=w_{o}=w_{pl}=1$.
-  • Derive impact rating: Based on the computed impact score of each threat, we derive the corresponding ratings using the thresholds defined in Table 2.
-This rating helps prioritise threats based on their potential severity and guides subsequent risk mitigation in our analysis.
-    Table 2: Mapping of computed impact scores to ratings
-    Impact Score Rating
+  • Assess impact level: By definition in the HEAVENS framework, the impact level of a specific threat is an estimate of the expected loss for various stakeholders when the threat is realised.
+To estimate the impact, we use four parameters directly related to the security objectives defined in HEAVENS: safety (s), financial (f), operational (o), and privacy (p).
+    • Safety Impact: Refers to the safety of vehicle occupants, road users, and infrastructure.
+    • Financial Impact: Encompasses all direct and indirect financial damages incurred by stakeholders.
+    • Operational Impact: Pertains to operational damages that have minimal or no safety or financial consequences. For example, the loss of secondary functionalities such as entertainment systems.
+    • Privacy Impact: Deals with damages resulting from privacy violations.
+    Each of the four parameters is assigned one of four impact levels: None, Low, Medium, or High.
+For the purposes of this experience report, we used the impact level definitions from [12].
+Impact levels are typically assigned by automotive domain experts who have detailed knowledge of the automotive system, its vulnerabilities, and corresponding impacts on security objectives.
+In our running example, we assign impact levels based on our assessment of potential safety, operational, financial, and privacy consequences.
+This assessment is based on our understanding from the literature, and may differ from evaluations of experts in the field.
+    Next, numerical values are assigned to the impact levels.
+These values are chosen in a logarithmic scale (0, 1, 10, 100) to effectively capture the exponential increase in the impact levels.
+This helps in clearly differentiating between the levels of impact related to the security objectives, and ensures that high-impact threats are given appropriate priority.
+    Table 10 [8] categorises the impact levels and parameter values according to the security objectives.
+  • Calculate impact score: After estimating the individual impact levels of the security objectives and their corresponding parameter values, the overall impact score is computed using the following equation:
+    $\small I=w_{s}i_{s}+w_{o}i_{o}+w_{f}i_{f}+w_{p}i_{p}$ (1)
+    Normalisation is necessary to ensure that the overall impact score is on a consistent scale, allowing for fair and meaningful comparisons across different systems and contexts.
+This uniformity is crucial when comparing the cybersecurity posture of various systems, such as between automotive systems.
+    The overall normalised impact score is computed using the following equation:
+    $I_{\text{norm}}=\frac{w_{s}i_{s}+w_{o}i_{o}+w_{f}i_{f}+w_{p}i_{p}}{100*(w_{s}+w_{o}+w_{f}+w_{p})}$ (2)
+    Weight values $w_{s}$, $w_{o}$, $w_{f}$, and $w_{p}$ are selected based on the specific needs of the use case being evaluated.
+In the case of an automotive system, the safety impact is the most severe consequence for stakeholders, such as vehicle occupants, road users, and infrastructure.
+Comparatively, financial, operational, and privacy impacts may have lower consequences.
+As such, the weights are selected as: $w_{s}=10$ and $w_{o}=w_{f}=w_{p}=1$.
+Equations 1 and 2 are simplified as follows:
+    $I=10i_{s}+i_{o}+i_{f}+i_{p}$ (3)
+    $\small I_{\text{norm}}=\frac{10i_{s}+i_{o}+i_{f}+i_{p}}{1300}$ (4)
+  • Assign impact rating: Each threat is assigned an impact rating based on the impact score $I_{\text{norm}}$ of each assessed threat, matched against thresholds tabulated in Table 2, as referenced from [12].
+    Table 2: Impact rating table
+    Impact Score Impact Rating
     0.00 None
@@
     $1.000.45\leq x<$ 1.00 Critical
-The final impact evaluation is tabulated in Figure 4.
-Figure 4: Impact assessment based on HEAVENS 2.0
-Step 5: Model attacks
+For illustration, we show how the impact rating of Threat ID 9146 from Table 9 [8] is derived using this methodology.
+In this specific threat, an adversary may reverse engineer the head unit firmware to find sensitive information, resulting in information disclosure from the IVI System’s local maps data store.
+Using the definitions of the impact assessment table (Table 10 [8]), we evaluate $i_{s}=0$, $i_{o}=0$, and $i_{f}=0$, as there are no injuries or discernible operational and financial effects resulting from the disclosure of information.
+However, disclosure of personally identifiable information or passwords from the data store can result in privacy violations leading to abuses such as impersonation of the victim to perform actions with stolen identities.
+As such, the impact level for privacy is assessed as Medium and $i_{p}=10$.
+Using Equation 2, $I_{\text{norm}}\ =0.0077$, which corresponds to a Low impact rating.
+Table 3: Impact rating example
+Threat ID $i_{s}$ $i_{o}$ $i_{f}$ $i_{p}$ Impact Score Impact Rating
+9146 0 0 0 10 0.0077 Low
+4.5 Step 5: Model attacks
 In this step, we use an attack tree to model attacker goals and map out attack paths associated with threat and attack vectors identified earlier using the Microsoft Threat Modelling Tool.
@@
 The attack tree depicted in Figure 3 illustrates the potential attack paths these two personas could employ to achieve their respective goals.
-Figure 5: Attack tree for IVI system
-Step 6: Evaluate attack feasibility
-Next, we use CVSS version 3.1 to analyse and compute the attack feasibility ratings.
-CVSS provides a standardised method for assessing the severity of vulnerabilities based on their impact and exploitability factors.
-The CVSS 3.1 base score is computed using Equation 2.
-Base Score=Round up(0.6×Impact+0.4×Exploitability1.5Impact $\tiny\text{Base Score}=\text{Round up}\left(\frac{{0.6\times\text{Impact}+0.4\times\text{Exploitability}-1.5}}{{\text{Impact Factor}+0.8}}\right)$ (2)
-where Impact, Exploitability, and Impact Factor are derived from the following components:
-  • Impact Metrics (C, I, A):
-    • Confidentiality (C): The potential impact on confidentiality.
-    • Integrity (I): The potential impact on integrity.
-    • Availability (A): The potential impact on availability.
-  • Exploitability Metrics (AV, AC, PR, UI):
+Figure 3: Attack tree for In-Vehicle Infotainment (IVI) system
+4.6 Step 6: Evaluate attack feasibility
+Next, we use CVSS version 3.1 to analyse and compute attack feasibility.
+CVSS provides a standardised method for assessing the severity of vulnerabilities based on their impact and exploitability factors, as described below.
+  • Impact Metrics:
+    • Confidentiality (C): Impact on the confidentiality of data, such as driver information, vehicle location, and system diagnostics.
+    • Integrity (I): Impact on the integrity of automotive systems, including the accuracy and reliability of data and control signals within the vehicle.
+    • Availability (A): Impact on the availability of automotive functions and services, such as the ability to start the vehicle or use navigation systems.
+  • Exploitability Metrics:
     • Attack Vector (AV): How the vulnerability is accessed.
@@
 If the vulnerability affects other system components, the Scope is considered Changed (S), otherwise, it is Unchanged (U).
-Additionally, the CVSS temporal score takes into account factors such as exploit code maturity, remediation level, and report confidence. It is calculated using Equation 3.
-Temporal Score=Round up(Base Score×Exploit Code Maturity×Remediation Level×Report $\tiny\begin{split}\text{Temporal Score}=\text{Round up}(\text{Base Score}\times\text{Exploit Code Maturity}\times\\ \text{Remediation Level}\times\text{Report Confidence})\end{split}$ (3)
-Figure 6: Attack feasibility ratings based on CVSS version 3.1
-Figure 6 lists the attack feasibility ratings computed using the CVSS version 3.1 calculator (https://nvd.nist.gov/vuln-metrics/cvss/v3-calculator).
-These ratings are based on the information available at the time of assessment during the early design stage of the IVI system.
-Figure 7: Risk ratings evaluated using a 5x5 risk matrix
-Step 7: Evaluate risks
-In this step, we evaluate the cybersecurity risks of the IVI system using a 5x5 risk matrix. The corresponding risk ratings for identified threats are tabulated in Figure 7.
+In this experience report, instead of using the CVSS base score, which is common in the literature, we use the CVSS temporal score to incorporate and emphasise the temporal aspects of the vulnerability, such as exploit code maturity, remediation level, and report confidence.
+These are shown at Equations 5 and 6.
+CVSS Base Score=Round up(0.6×Impact+0.4×Exploitability1.5Impact $\text{CVSS Base Score}=\text{Round up}\left(\frac{{0.6\times\text{Impact}+0.4\times\text{Exploitability}-1.5}}{{\text{Impact Factor}+0.8}}\right)$ (5)
+CVSS Temporal Score=Round up(CVSS Base Score×Exploit Code Maturity×Remediation Level×Report $\begin{split}\text{CVSS Temporal Score}=\text{Round up}(\text{CVSS Base Score}\times\text{Exploit Code Maturity}\times\\ \text{Remediation Level}\times\text{Report Confidence})\end{split}$ (6)
+Next, the CVSS temporal scores for the identified threats are computed using the CVSS version 3.1 calculator (https://nvd.nist.gov/vuln-metrics/cvss/v3-calculator), then mapped to qualitative severity ratings using Table 4 as referenced from the CVSS 3.1 Specification Document.
+The severity ratings, which we refer to as attack feasibility in this experience report, are tabulated in Table 5.
+Table 4: Qualitative severity rating scale
+CVSS Score Severity Rating (Attack Feasibility)
+0.0 - 3.9 Low
+4.0 - 6.9 Medium
+7.0 - 8.9 High
+9.0 - 10.0 Critical
+Table 5: Attack feasibility based on CVSS version 3.1
+Threat ID AV AC PR UI S C I A E RL RC CVSS Score Attack Feasibility
+9158 A L N N C N L H P U U 7.1 High
+9132 A H N N C N H H U U U 6.7 Medium
+9089 P L N R U H H H U U U 5.6 Medium
+9144, 9060 A L L R U H H H U U U 6.2 Medium
+9146 P H N R U H L N U U U 3.9 Low
+4.7 Step 7: Evaluate risks
+In this step, we evaluate the cybersecurity risks of the IVI system using a 5x5 risk matrix illustrated in Figure 4.
+In the context of automotive impact assessment, where the stakes can be high and risks multifaceted, a 5x5 risk matrix provides the depth and granularity needed to accurately assess and manage risks, compared to a simpler 3x3 risk matrix.
+The risk ratings for the identified threats are tabulated in Table 6.
+Figure 4: 5x5 risk matrix
+Table 6: Risk ratings based on attack feasibility and impact ratings
+Threat ID Description Attack Method Source Attack Feasibility Impact Rating Risk Rating
+9132 Cause a denial-of-service of the central gateway, affecting vehicle functions Inject or flood invalid data to the central gateway causing it to crash or stop. IVI System High Critical Extreme
+9159 Cause gateway to misbehave, crash or stop, affecting vehicle functions Exploit vulnerabilities, through outdated software or configuration weaknesses. IVI System Medium High High
+9089 Elevate privileges on the IVI System (via USB) Exploit vulnerabilities, through outdated software or configuration weaknesses, via USB removable media. USB Medium Low Medium
+9144, 9060 Elevate privileges on the IVI System (via Bluetooth or Wi-Fi) Exploit vulnerabilities, through outdated software or configuration weaknesses, via Bluetooth or Wi-Fi.
+Bluetooth or Wi-Fi Medium Low Medium
+9146 Access sensitive data e.g.
+passwords from IVI System Embedded Data Store Reverse engineer the head unit firmware to find keys to access the IVI System Embedded Data Store.
+Physical Access Low Low Medium
 These risk ratings serve as a guideline for the Original Equipment Manufacturer (OEM) of the vehicle or the manufacturer of the IVI system.
-They can use these ratings to prioritise the implementation of mitigation controls during the design phase of the system life cycle.
+They can use these ratings to prioritise the implementation of mitigation controls throughout the system life cycle.
 This approach ensures that adequate attention and resources are allocated to address the most critical cybersecurity concerns identified through this comprehensive analysis.
-Step 8: Review, update, and iterate
+4.8 Step 8: Review, update, and iterate
 In this final step, we continuously review the cybersecurity analysis and update the risk assessment based on new threat information, insights or changes in the automotive system environment.
@@
 This iterative cycle of review and improvement is crucial for maintaining the automotive system’s long-term cybersecurity posture.
-Updating the Security Model After a Vulnerability Disclosure
-In the earlier section, we discussed the TARA framework as part of our approach to security modelling for automotive systems.
-Generally, this is performed during the early stages of system life cycles.
-However, evolving cybersecurity threats necessitate the incorporation of new knowledge gained from ongoing security monitoring and threat intelligence throughout the system’s life cycle.
-ACTISM can be seamlessly integrated with the TARA framework to create a threat-informed dynamic security modelling methodology.
+5 Updating the Security Model After a Vulnerability Disclosure
+ACTISM enhances security modelling by introducing a dynamic and iterative methodology, resulting in a threat-informed, evolving security model.
+This approach addresses evolving cybersecurity threats that necessitate the incorporation of new knowledge gained from security monitoring and threat intelligence throughout the system’s life cycle.
 We now apply this methodology to a real-world example using the Tesla Electric Vehicle (EV) jail-break vulnerability disclosure [14].
+Although this vulnerability is specific to the Tesla EV platform, ACTISM is designed to be applicable to all automotive systems (including non-EVs) and their relevant vulnerabilities.
 After gathering technical details from the vulnerability disclosure, we review the IVI system security model developed in the earlier section and update the CVSS metrics as necessary.
@@
   • Report Confidence (RC) From Unknown (U) to Confirmed (C), in view of the disclosure from the security researchers.
-The updated risk rating based on these revised metrics is shown in Figure 8. In summary, the corresponding risk changed from Medium to High.
+With these updated parameters, the attack feasibility increased from 2 (Low) to 3 (Medium), as shown in Table 7.
+Table 7: Updated attack feasibility following vulnerability disclosure
+Threat ID AV AC PR UI S C I A E RL RC CVSS Score Attack Feasibility
+9158 A L N N C N L H P U U 7.1 High
+9132 A H N N C N H H U U U 6.7 Medium
+9089 P L N R U H H H U U U 5.6 Medium
+9144, 9060 A L L R U H H H U U U 6.2 Medium
+9146 P H N R C H H H P T C 6.4 Medium
+Besides attack feasibility, the impact assessment of Threat ID 9146 in the security model also needs updating.
+Using the definitions from the impact assessment table (Table 10 [8]), we estimate the impact on safety to increase from None to Low ($i_{s}=10$), as there is now a possibility of light and moderate injuries, should the attacker access and manipulate certain vehicle functions via the vulnerability.
+We also expect the operational impact to increase from None to Low ($i_{o}=10$), as there may be a degradation of primary vehicle functions.
+The financial impact is also expected to increase from None to Low ($i_{f}=10$), as the financial damage from injury claims may be substantial but not threatening the existence of the manufacturer.
+Lastly, the privacy impact is expected to increase from Low to High ($i_{p}=100$), as the privacy violation of multiple vehicle owners and drivers may lead to extensive media coverage and severe loss of market share.
+Using Equation 2 to compute the updated normalised impact score, $I_{\text{norm}}\ =0.1692$, which corresponds to a High impact rating (Table 8).
+Table 8: Updated impact rating following vulnerability disclosure
+Threat ID $i_{s}$ $i_{o}$ $i_{f}$ $i_{p}$ Updated Impact Score Updated Impact Rating
+9146 10 10 10 100 0.1692 High
+With the updated attack feasibility (Medium) and impact rating (High), the overall risk rating for Threat ID 9146 increased from Medium to High, in accordance with the 5x5 risk matrix (Figure 4).
 This ability to perform timely updates to the risk assessment is crucial for maintaining robust cybersecurity in automotive systems.
-By continuously adapting to new threats and changes in attack feasibility and impact, proactive measures can be implemented to safeguard automotive system safety and functionality.
-This dynamic approach not only enhances the resilience of automotive systems against evolving cybersecurity threats but also reinforces the importance of informed decision-making in the face of emerging risks.
-Ultimately, it empowers stakeholders to take appropriate actions that protect both the automotive system and its users.
-Figure 8: Updated risk rating following vulnerability disclosure
-Discussion
-Unlike early-stage security modelling, where OEMs and IVI system manufacturers implement comprehensive measures, addressing cyber-physical vulnerabilities later in the life cycle often involves temporary fixes or workarounds.
-These may not fully mitigate the risks, so vehicle owners can use risk ratings to make informed decisions.
-By adopting additional precautions, they can enhance the cybersecurity resilience of their vehicles throughout their life cycle.
-This underscores the need for a continuous, iterative, and threat-informed security modelling process.
-Ongoing threat intelligence gathering, including real-time monitoring and tracking vulnerability disclosures, is crucial.
-A comprehensive repository of vulnerability attributes, stored in a standardised format such as JSON, enables automation.
-Tools can parse this data, extract relevant information, and enrich the security model via workflows like TARA.
-CVSS scores and risk ratings could then be dynamically recalculated based on updated threat context.
-Security models can be formalised using Domain-Specific Languages (DSLs) such as the Meta Attack Language (MAL), allowing clearer descriptions of threats and vulnerabilities.
-Furthermore, adversary tactics from the MITRE ATT&CK knowledge base could be incorporated to enhance these models.
-Machine learning approaches could also help prioritise vulnerabilities by analysing their impact, using historical data and threat intelligence to improve automotive cybersecurity.
-Conclusion
-Automotive systems are increasingly facing cybersecurity threats, with attackers targeting vulnerabilities that pose risks to data, functionality, and safety.
-Ensuring the security of automotive systems is critical to maintaining their integrity and reliability.
-We introduced ACTISM (Automotive Consequence-Driven and Threat-Informed Security Modelling), a novel approach designed to enhance the resilience of automotive systems by dynamically updating their cybersecurity posture.
-ACTISM responds to evolving threats and attacker tactics, ensuring security models remain relevant throughout a system’s lifecycle.
-Unlike traditional models, ACTISM continuously integrates new data from threat intelligence and security monitoring, keeping protection up to date.
-It also captures the multi-agent nature of cyber-physical attacks, making it applicable to automotive domains.
-This integration strengthens the accuracy and completeness of security models, reflecting the complexity of real-world cyber-physical attacks.
-Cybersecurity incidents in automotive systems differ from traditional IT breaches due to their potential physical consequences, such as vehicle misnavigation or compromised safety.
-ACTISM’s consequence-driven approach ensures that cyber-physical risks are considered, directly addressing the operational and safety integrity of automotive systems.
-To demonstrate ACTISM’s practical application, we developed a comprehensive security model for an IVI system.
-This involved identifying key assets, conducting threat modelling, and performing risk assessments, resulting in a robust security framework.
-A real-world case study further highlighted ACTISM’s ability to adapt to emerging threats and its potential for broader use in other automotive technologies.
-In conclusion, ACTISM offers a proactive, adaptable approach to automotive cybersecurity, addressing current challenges while preparing for the evolving landscape of automotive technology.
+Ultimately, it empowers stakeholders to adopt additional precautions and improve the cybersecurity resilience of their vehicles throughout their life cycle.
+6 Future directions
+While ACTISM provides an approach to dynamically updating security models with both cyber and physical threats, it is not without its limitations.
+One of the primary challenges is how the accuracy of the model can be validated given the constantly evolving nature of threats and the potential for incomplete or inaccurate threat intelligence data.
+Moreover, while ACTISM’s adaptability to new threats is a strength, it can also pose a challenge in maintaining the models’ integrity and consistency over time.
+Despite these limitations, the potential future directions outlined below can address these challenges and enhance the robustness of ACTISM.
+Ongoing threat intelligence gathering, including real-time monitoring and tracking of vulnerability disclosures, is crucial for keeping the security model current.
+Establishing a comprehensive repository of vulnerability attributes, stored in a standardised format such as XML or JSON, can enable automation and streamline the updating process.
+Additionally, formalising security models using Domain-Specific Languages (DSLs) like Automation Markup Language (AutomationML) and Meta Attack Language (MAL) can provide clearer descriptions of threats and vulnerabilities.
+Incorporating adversary tactics from the MITRE ATT&CK knowledge base can further enhance these models by offering insights into real-world attack strategies.
+Moreover, leveraging machine learning approaches to prioritise vulnerabilities based on their impact, utilising historical data and threat intelligence, can significantly improve automotive cybersecurity.
+These potential future directions can ensure that automotive cybersecurity remains resilient and adaptive to emerging cybersecurity challenges.
+7 Conclusion
+The cybersecurity landscape for automotive systems is rapidly evolving, with attackers increasingly targeting vulnerabilities that threaten data integrity, functionality, and safety.
+Traditional static security assessment methodologies, such as STRIDE, often fall short in addressing the dynamic and multi-faceted nature of these threats.
+There is a clear knowledge gap in the ability to model and respond to security threats in a dynamic, iterative manner tailored specifically for automotive systems.
+In response, we introduced ACTISM (Automotive Consequence-Driven and Threat-Informed Security Modelling), an integrated approach designed to enhance the resilience of automotive systems by dynamically updating their cybersecurity posture.
+ACTISM effectively responds to evolving cyber-physical threats and attacker tactics, ensuring that security models remain robust and relevant throughout a system’s lifecycle.
+To illustrate the practical application of ACTISM, we developed a comprehensive security model for an IVI system.
+This process included identifying key assets, conducting thorough threat modelling, and performing risk assessments, ultimately resulting in a robust security framework.
+A real-world case study involving Tesla’s IVI system further demonstrated ACTISM’s ability to adapt to emerging threats and underscored its potential for broader applicability across various automotive technologies.
+In conclusion, ACTISM addresses the critical knowledge gap in automotive cybersecurity by providing a proactive, adaptable approach to security modelling.
+This approach not only tackles current challenges but also prepares for the future landscape of automotive technology.

### v3

@@
 Abstract
-Cybersecurity threats in complex cyber-physical systems pose significant risks to system functionality and safety.
-This experience report introduces ACTISM (Automotive Consequence-Driven and Threat-Informed Security Modelling) — an integrated security modeling approach that enhances the resilience of automotive systems by dynamically updating their cybersecurity posture in response to prevailing threats, attacker tactics, and their impact on system functionality and safety.
+Evolving cybersecurity threats in complex cyber-physical systems pose significant risks to system functionality and safety.
+This experience report introduces ACTISM (Automotive Consequence-Driven and Threat-Informed Security Modelling), an integrated security modelling framework that enhances the resilience of automotive systems by dynamically updating their cybersecurity posture in response to prevailing and evolving threats, attacker tactics, and their impact on system functionality and safety.
 ACTISM addresses the existing knowledge gap in static security assessment methodologies by providing a dynamic and iterative framework.
 We demonstrate the effectiveness of ACTISM by applying it to a real-world example of the Tesla Electric Vehicle’s In-Vehicle Infotainment system, illustrating how the security model can be adapted as new threats emerge.
-We also outline avenues for future research and development in this area, including automated vulnerability management workflows for automotive systems.
+We also report the results of a practitioners’ survey on the usefulness of ACTISM and its future directions.
+The survey highlights avenues for future research and development in this area, including automated vulnerability management workflows for automotive systems.
 Keywords:
@@
 1 Introduction
-Automotive systems have transformed into complex cyber-physical systems brimming with cutting-edge features.
-In contrast to their predecessors, today’s automotive systems are equipped with a variety of advanced sensors and connectivity options, such as Control Area Networks (CAN), Local Interconnect Networks (LIN), FlexRay, and Ethernet.
-They also offer external connectivity through Bluetooth, Wi-Fi, and GPS.
-In addition, the integration of feature-rich In-Vehicle Infotainment (IVI) systems and powerful computing units for autonomous driving continues to push the boundaries of what is possible in automotive technology.
-However, this technological evolution also broadens the attack surface, making automotive systems more susceptible to cybersecurity attacks.
-Cybersecurity threats and vulnerabilities in the automotive industry are well documented, and attackers exploit these vulnerabilities to compromise vehicle security, access sensitive data, and even take control of vehicle functions.
-Recent literature has increasingly highlighted the challenges of defending against cybersecurity threats in automotive systems.
-In their survey on security attacks and defence techniques for Connected and Autonomous Vehicles (CAVs), Pham and Xiong [11] observed that attackers could potentially compromise one vehicle component after another.
-To mitigate this risk, they emphasised the need for manufacturers to strengthen the security of all system components to ensure the overall protection of CAVs.
-While this approach may be feasible during the early stages of an automotive system’s life cycle, evolving cybersecurity threats are likely to render it less effective over time.
-Given the rapid increase in both threats and the complexity of automotive systems, it is unrealistic to expect that an automotive system’s initial cybersecurity features will counter all emerging risks throughout its entire life cycle.
-This is where threat-informed dynamic threat and attack models become crucial, offering a proactive and adaptive approach to designing secure automotive systems.
-Dynamic threat and attack models provide engineers with a structured framework to analyse and mitigate cybersecurity risks, ensuring that automotive systems remain protected against evolving cybersecurity threats.
-Additionally, they enable consumers to make informed decisions when selecting vehicles based on their cybersecurity robustness.
-However, many of the threat and attack models discussed in the literature do not fully address the unique cybersecurity requirements of modern automotive systems. These requirements include:
+Automotive systems have evolved into sophisticated cyber-physical systems, packed with cutting-edge technologies.
+Unlike their predecessors, today’s vehicles are equipped with advanced sensors, including LiDAR (Light Detection and Ranging), radar, and cameras, alongside a variety of connectivity options such as Control Area Networks (CAN), Local Interconnect Networks (LIN), FlexRay, and Ethernet.
+External connectivity is further enhanced through Bluetooth, Wi-Fi, and GPS, enabling seamless integration with external devices and services.
+Additionally, the integration of feature-rich In-Vehicle Infotainment (IVI) systems and powerful computing units for autonomous driving continues to push the boundaries of automotive innovation.
+However, this technological evolution significantly broadens the attack surface, making automotive systems increasingly vulnerable to cybersecurity threats.
+Cybersecurity threats and vulnerabilities in the automotive industry are well-documented, with attackers exploiting these weaknesses to compromise vehicle security, access sensitive data, and even take control of critical functions.
+Recent literature underscores the growing challenges of defending automotive systems against such threats.
+For instance, Pham and Xiong [13], in their survey on security attacks and defence techniques for Connected and Autonomous Vehicles (CAVs), highlighted how attackers can sequentially compromise vehicle components.
+They stressed the importance of securing all system elements to ensure comprehensive protection.
+While this approach is viable during the early stages of a vehicle’s life cycle, it becomes less effective over time as cybersecurity threats evolve in both complexity and scale.
+The rapid advancement of automotive technologies further exacerbates this issue, making it impractical to rely solely on initial cybersecurity measures to address emerging risks throughout a vehicle’s operational life.
+This is where dynamic threat and attack models prove invaluable.
+By providing a proactive and adaptive framework, these models enable engineers to continuously analyse and mitigate cybersecurity risks, ensuring systems remain resilient against evolving threats.
+Additionally, they empower consumers to make informed decisions by evaluating a vehicle’s cybersecurity robustness.
+However, many existing threat and attack models do not fully address the unique cybersecurity requirements of modern automotive systems.
+These requirements include:
   • Interconnectivity: Modern vehicles are highly interconnected systems, comprising numerous Electronic Control Units (ECUs) that communicate with each other.
@@
 For instance, threat models developed using Microsoft’s STRIDE (Spoofing, Tampering, Repudiation, Information Disclosure, Denial of Service, and Elevation of Privilege) primarily focus on data flows and information security.
-However, STRIDE was not designed with automotive systems in mind, making its application in this context somewhat limited.
-Consequently, STRIDE often overlooks critical aspects of how cybersecurity breaches can directly affect automotive functions and safety.
-As a result, the reliability of the general outcomes from STRIDE is questionable when applied to the unique characteristics of automotive environments.
-In other words, threat modelling of automotive systems often does not consider the consequences of cybersecurity attacks beyond information security.
-Similarly, attack models designed for traditional IT systems do not fully align with the multi-path tactics and techniques employed by attackers seeking to gain control of a vehicle.
-These techniques often involve a combination of physical and cyber actions, necessitating a more comprehensive and dynamic approach to security modelling.
-These findings highlight the need to tailor security modeling approaches for automotive systems so that they can adequately capture the unique cybersecurity threats faced in these environments.
+As a result, current threat modelling practices for automotive systems often fail to consider the broader consequences of cybersecurity attacks beyond information security [5].
+They often overlook critical aspects of how cybersecurity breaches can directly impact automotive functions and safety, raising questions about the reliability of their outcomes when applied to the unique characteristics of automotive environments.
+Similarly, attack models designed for traditional IT systems do not fully account for the multi-path tactics and techniques attackers use to gain control of a vehicle [2][6].
+These techniques often involve a combination of physical and cyber actions, highlighting the need for a more comprehensive and dynamic approach to security modelling.
+These findings underscore the importance of tailoring security modelling approaches to automotive systems, ensuring they adequately capture the unique cybersecurity threats these environments face.
 Contributions.
-In this experience report, we introduce ACTISM (Automotive Consequence-Driven and Threat-Informed Security Modelling) — an integrated security modeling approach that enhances the resilience of automotive systems by dynamically updating their cybersecurity posture in response to prevailing threats, attacker tactics, and their impact on system functionality and safety.
+In this experience report, we introduce ACTISM (Automotive Consequence-Driven and Threat-Informed Security Modelling) — an integrated security modelling framework that enhances the resilience of automotive systems by dynamically updating their cybersecurity posture in response to prevailing and evolving threats, attacker tactics, and their impact on system functionality and safety.
 This approach addresses the existing knowledge gap in static security assessment methodologies by providing a dynamic and iterative framework.
-We demonstrate the effectiveness of ACTISM by applying it to an IVI system, illustrating how the model can be adapted as new threats emerge.
-Unlike existing security modelling approaches, ACTISM integrates various frameworks to dynamically update security models with new information, acquired through threat intelligence and ongoing security monitoring.
-This keeps security models relevant against evolving attacker tactics, techniques and procedures (TTPs) even if the models were developed in early stages of system development.
-ACTISM also allows practitioners to apply the security models in both IT and automotive domains by using attack paths to capture the multi-path and multi-agent aspects of real-world cyber-physical attacks.
-In order to bolster the completeness of the security model, ACTISM integrates the following threat analysis and risk assessment frameworks:
-  • Microsoft’s STRIDE methodology.
-  • HEAVENS (HEAling Vulnerabilities to ENhance Software Security and Safety), a structured and model-based hazard analysis and risk assessment method specifically designed for automotive systems.
-  • Attack Tree, a hierarchical diagram representing potential attack paths, illustrating how an attacker could exploit system vulnerabilities to achieve specific goals.
+We demonstrate the effectiveness of ACTISM by applying it to an IVI system, showing how the model can be adapted as new threats emerge.
+Unlike existing security modelling frameworks, ACTISM integrates multiple methodologies to dynamically update security models with new information gathered through threat intelligence and continuous security monitoring.
+This ensures that the models remain relevant against evolving attacker tactics, techniques, and procedures (TTPs), even if they were initially developed during the early stages of system design.
+ACTISM also enables practitioners to apply these security models across both IT and automotive domains by utilising attack paths to accurately represent the multi-path and multi-agent nature of real-world cyber-physical attacks.
+To enhance the comprehensiveness of its security model, ACTISM applies and adapts the following threat analysis and risk assessment frameworks:
+  • STRIDE, a widely-used threat modelling framework developed by Microsoft to characterize security threats in software systems.
+The acronym STRIDE stands for six types of threats: Spoofing, Tampering, Repudiation, Information Disclosure, Denial of Service, and Elevation of Privilege.
+In addition to information security, ACTISM reports how physical threats can be modeled using STRIDE in conjunction with other frameworks (Section 4).
+  • HEAVENS (i.e., HEAling Vulnerabilities to ENhance Software Security and Safety), a structured and model-based hazard analysis and risk assessment method specifically designed for automotive systems.
+  • Attack Tree, a hierarchical diagram representing potential attack paths, used for modeling how an attacker could exploit system vulnerabilities to achieve specific goals.
   • Common Vulnerability Scoring System (CVSS), an open industry standard for assessing the severity of computer system security vulnerabilities.
-CVSS (Common Vulnerability Scoring System): This model is also static as it provides a standardised method for evaluating the severity of vulnerabilities at a particular point in time.
-It does not dynamically adapt to new threats without manual intervention.
-  • Threat Analysis and Risk Assessment (TARA) methodology from the ISO/SAE 21434:2021 ("Road Vehicles - Cybersecurity Engineering") Standard, jointly developed by the International Organisation for Standardisation (ISO) and the Society of Automotive Engineers (SAE), for managing cybersecurity threats and vulnerabilities throughout the life cycle of a vehicle.
-Table 1 lists the differences between ACTISM and these security modelling frameworks, and their respective limitations.
-In contrast with other security modelling frameworks, the ACTISM framework integrates various frameworks to be dynamic, continuously adapting to new threats and updating the automotive system’s cybersecurity posture.
-This distinction sets it apart from the static models, providing a more robust approach to maintaining automotive cybersecurity.
-Table 1: Comparison of ACTISM with other security modelling frameworks
-Framework Type Limitations
-ACTISM Dynamic Relies on the integration of various security modelling frameworks.
-STRIDE Static Limited to specific threat categories. Does not dynamically adapt to new threats without manual intervention.
-Attack Trees Static Requires manual updates and may not cover all emerging threats. Does not dynamically adapt to new threats without manual intervention.
-HEAVENS Static Primarily focuses on high-level security analysis which may not cover detailed, system-specific vulnerabilities and threats.
-Does not dynamically adapt to new threats without manual intervention.
-CVSS Static Does not dynamically adapt to new threats without manual intervention.
-TARA Static Does not dynamically adapt to new threats without manual intervention.
-Although frameworks such as STRIDE and CVSS were designed for IT security threats, they can also be applied to physical threats.
-As an example, tampering with physical components can lead to unauthorised access, data manipulation, or service disruption.
-As an integrative approach, ACTISM provides guidance on how both cyber and physical attacks can be modeled using these frameworks to capture the various entry points and methods attackers might use, ensuring comprehensive security coverage across both cyber and physical domains.
-To demonstrate the practical application of ACTISM, we developed a comprehensive security model for an IVI system.
-We outline the step-by-step process of building the security model, emphasising the integration of threat intelligence and contextual information to accurately characterise and assess cybersecurity risks.
-Furthermore, we demonstrate the utility of ACTISM by illustrating its adaptive response to a real-world vulnerability disclosure affecting the IVI system.
-This case study underscored ACTISM’s effectiveness in addressing emerging threats and evolving security challenges in automotive cybersecurity.
-Furthermore, the study showcased the versatility of ACTISM, demonstrating its potential applicability to other automotive systems beyond the IVI, making it a robust tool for safeguarding a wide range of vehicle technologies.
-Finally, we outline avenues for future research and development in this area, including automated vulnerability management workflows.
+This model is typically used in a static manner, i.e., the severity scores are usually not updated for new or evolving threats.
+ACTISM shows how CVSS can be applied dynamically (Section 5).
+  • Threat Analysis and Risk Assessment (TARA) methodology adopted from the ISO/SAE 21434:2021 ("Road Vehicles - Cybersecurity Engineering") Standard, jointly developed by the International Organisation for Standardisation (ISO) and the Society of Automotive Engineers (SAE), for managing cybersecurity threats and vulnerabilities throughout the vehicle’s life cycle.
+While frameworks like STRIDE and CVSS were originally designed for IT security, they can also be applied to physical threats.
+For instance, tampering with physical components can lead to unauthorised access, data manipulation, or service disruption.
+ACTISM takes an integrative approach, combining these frameworks to model both cyber and physical attacks.
+This ensures comprehensive coverage of potential entry points and attack methods, addressing the multi-faceted nature of modern automotive threats.
+To demonstrate ACTISM’s practical application, we developed a detailed security model for an In-Vehicle Infotainment (IVI) system.
+The step-by-step process highlights the integration of threat intelligence and contextual information to accurately characterise and assess cybersecurity risks.
+Additionally, we illustrate ACTISM’s adaptive capabilities by simulating its response to a real-world vulnerability disclosure affecting the IVI system.
+This case study underscores ACTISM’s effectiveness in addressing emerging threats and evolving security challenges, while also reporting its versatility for application across other automotive systems beyond IVI.
+Finally, we report the results of a practitioners’ survey on the usefulness of ACTISM and its future directions.
+The survey highlights avenues for future research and development in this area, including automated vulnerability management workflows for automotive systems.
+These advancements aim to further enhance ACTISM’s ability to safeguard automotive systems against increasingly sophisticated cybersecurity threats.
 2 Related Work
-Various works have attempted to address security modelling in automotive systems.
 Das et al.
@@
 Secondly, the CVSS version 2.0 framework adopted in Abuabed et al.
-[1] is no longer current.
-The newer versions of CVSS include improvements that deliver more granularity and accuracy in the resulting risk assessment.
+[1] lacks the granularity and accuracy that newer versions of CVSS can provide, which impacts the resulting risk assessment.
 Lastly, the framework does not adequately address threats in later stages of the life cycle.
 As automotive system life cycles are typically 10 to 20 years or longer, the vehicle system could become increasingly vulnerable to new types of attacks that were not anticipated during the initial security assessment.
+Other literature relevant to automotive cybersecurity includes a study by Girdhar et al.
+[7] who used the STRIDE methodology to create a threat model focused on post-accident cyberattack event analysis for CAV.
 Chlup et al.
@@
 However, the approach does not perform attack feasibility estimation and risk assessment, which are key steps in the TARA workflow.
-Other literature relevant to automotive cybersecurity includes a study by Girdhar et al.
-[5] who used the STRIDE methodology to create a threat model focused on post-accident cyberattack event analysis for CAV.
-Additionally, Dang et al.
+Dang et al.
 [3] used a STRIDE framework to analyse and classify cybersecurity threats in an autonomous vehicle architecture, including the infotainment system.
@@
 This means that attacks that involve access to the vehicle’s internal architecture from within the vehicle are excluded in the analysis.
-The current landscape of security modelling highlights several limitations.
+In summary, the current landscape of security modelling highlights several limitations.
 Firstly, there are differing assumptions regarding which vehicle assets fall within the trust boundaries of security models.
@@
 These gaps underscore the need for an iterative and cohesive framework to effectively address the evolving challenges in automotive cybersecurity.
-3 Threat-informed Dynamic Security Modelling with ACTISM
-We propose ACTISM - a methodology for modelling security threats and attack vectors in a dynamic, iterative manner, specifically designed for automotive systems.
-We will illustrate the application of ACTISM to an IVI system as a running example and apply ACTISM to a real-world case study, where we demonstrate its ability to dynamically adapt to new threats and improve the security of automotive systems.
-This case study provides concrete evidence of the framework’s applicability and benefits, distinguishing our research from previous work.
-Current security modelling frameworks offer valuable insights but lack adaptability to evolving cyber-physical threats.
-ACTISM addresses this gap by introducing a framework that continuously updates the cybersecurity posture of automotive systems in response to both cyber and physical threats, attack vectors, and system vulnerabilities.
-Unlike static models, ACTISM incorporates threat intelligence through continuous monitoring, which ensures that the security model remains adaptive and effective in mitigating emerging risks.
-This iterative approach is vital because threat and attack models created in the early stages of system life cycles can become outdated or irrelevant over time as new vulnerabilities emerge.
-This gap widens, particularly when cyber-physical attacks occur like those targeting keyless entry systems [15], breaching the trust boundary within security models.
-ACTISM does not replace current threat and attack modelling frameworks.
-Instead, it streamlines their adoption and utilisation by cybersecurity practitioners.
-By integrating various security modelling frameworks into an iterative approach, ACTISM enhances the practicality and efficacy of addressing evolving cybersecurity threats and attacker capabilities.
-ACTISM comprises eight key steps as shown in Figure 1.
+3 High-Level Overview of ACTISM
+We propose ACTISM, a framework for modelling security threats and attack vectors in a dynamic, iterative manner, specifically designed for automotive systems.
+Unlike static models, ACTISM incorporates threat intelligence through continuous monitoring, which ensures that the security model remains adaptive and effective in mitigating emerging risks, like those targeting keyless entry systems [16], breaching the trust boundary within security models.
 Figure 1: Overview of ACTISM
-To enhance the practical utility of ACTISM among cybersecurity practitioners, we adopt the STRIDE methodology using a STRIDE-per-element approach, where threats are assessed against each element in the Data Flow Diagram.
-We choose not to use the STRIDE-per-interaction approach, as research has shown it yields a statistically significant lower level of completeness compared to the STRIDE-per-element approach [13].
-Furthermore, we utilise HEAVENS version 2.0 [9] for impact assessment.
+ACTISM is not designed to replace existing threat and attack modelling frameworks.
+Rather, it simplifies their adoption and application for cybersecurity practitioners.
+By integrating multiple security modelling frameworks into a cohesive, iterative process, ACTISM significantly improves the practicality and effectiveness of addressing evolving cybersecurity threats and attacker capabilities.
+As illustrated in Figure 1, ACTISM consists of eight key steps.
+These steps integrate multiple other threat and risk assessment frameworks, such as STRIDE, HEAVENS, attack trees, CVSS, and TARA, as outlined in Section 1.
+The details of each step will be expanded upon in Section 4.
+To enhance the practical utility of ACTISM among cybersecurity practitioners, we utilise HEAVENS version 2.0 [11] for impact assessment.
 To our knowledge, HEAVENS is the only framework designed to be fully compliant with the latest automotive cybersecurity standards, such as ISO/SAE 21434 and UNECE Regulation 155.
@@
 This comprehensive approach ensures that ACTISM remains robust and adaptive to the dynamic nature of cybersecurity threats in automotive systems.
-4 Applying ACTISM to an In-Vehicle Infotainment (IVI) System
-To illustrate the practical application of ACTISM, it is essential to understand the IVI system and its role in automotive cybersecurity.
-Legacy head units offered basic features such as radio reception and CD playback, with limited connectivity.
-In contrast, modern IVI systems are advanced digital platforms that integrate various functionalities beyond audio and navigation.
-These systems serve as central hubs, providing connectivity to external networks, managing complex data streams, enabling access to vehicle diagnostics, and supporting interactive user interfaces for both drivers and passengers.
-While modern IVI systems enhance convenience and operational efficiency, they also expose vehicles to increased cybersecurity risks.
-These systems are vulnerable to traditional threats, such as remote exploits and data breaches, as well as emerging threats like cyber-physical attacks targeting vehicle functionalities.
-The complexity of modern IVI systems significantly expands their attack surface.
-They integrate multiple wireless communication protocols—Wi-Fi, Bluetooth, and cellular networks—and connect to external services and mobile devices, increasing potential vulnerabilities.
-Malicious actors can exploit weaknesses in wireless protocols or insecure network connections to gain unauthorised access.
-Furthermore, advanced functionalities like voice recognition and artificial intelligence add layers of complexity, creating additional attack vectors.
-A dynamic approach like ACTISM allows for iterative security modelling throughout the IVI system’s life cycle, ensuring comprehensive protection against evolving threats.
+4 Applying ACTISM to an IVI System
+To demonstrate ACTISM’s utility, we use an In-Vehicle Infotainment (IVI) system as a running example and apply it to a real-world case study.
+To appreciate the significance of this application, it is essential to understand the role of IVI systems in modern vehicles and their associated cybersecurity challenges.
+Legacy head units offered basic functionalities such as radio reception and CD playback, with minimal connectivity.
+In contrast, modern IVI systems are sophisticated digital platforms that integrate a wide range of features beyond audio and navigation.
+They act as central hubs, managing complex data streams, enabling access to vehicle diagnostics, and providing interactive user interfaces for drivers and passengers.
+Additionally, they connect to external networks, mobile devices, and cloud-based services, significantly enhancing convenience and operational efficiency.
+However, these advancements also introduce substantial cybersecurity risks.
+Modern IVI systems are vulnerable to traditional threats like remote exploits and data breaches, as well as emerging cyber-physical attacks that target vehicle functionalities.
+The integration of multiple wireless communication protocols such as Wi-Fi, Bluetooth, and cellular networks expands the attack surface, providing malicious actors with opportunities to exploit weaknesses in wireless protocols or insecure network connections.
+Advanced features like voice recognition and artificial intelligence further increase complexity, creating additional attack vectors.
+A dynamic framework like ACTISM is uniquely suited to address these challenges.
+By enabling iterative security modelling throughout the IVI system’s life cycle, ACTISM ensures comprehensive protection against evolving threats, safeguarding both functionality and safety in modern vehicles.
 4.1 Step 1: Identify assets
@@
 These efforts aim to ensure a comprehensive understanding and coverage of all relevant assets within the IVI system.
-In this experience report, we reference the IVI HPC system DFD from Das et al.
+In this experience report, we build on the IVI HPC system DFD from Das et al.
 (2024) [4] to guide our identification of IVI system assets.
 This initial asset identification, listed below, serves as a foundational step for conducting further detailed analysis and security modelling.
-It is important to note that this DFD is an example and may not encompass newer types of attacks, such as dirty road attacks that exploit the fact that autonomous vehicles rely on sensors and cameras to interpret their environment.
-Attackers may place objects or patterns on the road that confuse these sensors, leading to incorrect navigation decisions or even accidents.
+While the referenced DFD provides a comprehensive framework for understanding IVI system components and interactions, we extend its scope to include emerging threats, such as dirty road attacks [14].
+These attacks exploit the fact that autonomous and semi-autonomous vehicles rely heavily on sensors and cameras to interpret their environment.
+Attackers may place malicious objects, patterns, or signals on the road to confuse these sensors, leading to incorrect navigation decisions, system malfunctions, or even accidents.
+By incorporating such attack vectors, we aim to enhance the cyber-physical security modelling process and address the evolving threat landscape faced by modern automotives.
   • In-vehicle modules
@@
 4.2 Step 2: Construct Data Flow Diagrams
-In the next step, we construct and analyse DFDs to identify the vehicle trust boundary and potential threat and attack vectors.
+In the next step, we construct and analyse DFDs to identify the vehicle trust boundary and potential threats and attack vectors, taking into consideration physical aspects.
 This process uses the Microsoft Threat Modelling Tool (Version 7.3.31026.3) along with the Automotive Threat Modelling Template [12] developed by the NCC Group.
@@
 Figure 2: Data Flow Diagram for an In-Vehicle Infotainment (IVI) System
-4.3 Step 3: Model threats
+4.3 Step 3: Model threats (cyber & physical)
 Next, we use the Microsoft Threat Modelling Tool to generate threats based on the previously constructed Data Flow Diagram (DFD) for the IVI system.
 The tool categorises these threats according to the STRIDE threat model automatically, providing an initial classification into categories such as Spoofing, Tampering, Repudiation, Information Disclosure, Denial of Service, and Elevation of Privilege.
+In addition to cyber threats, we also model physical threats, such as unauthorised physical access to the IVI system or tampering with hardware components.
 While the tool categorises threats by default and attempts to provide threat descriptions based on its own threat catalogues and templates, the initial output may be somewhat generic and not specifically tailored to the nuances of the IVI system’s operating environment.
@@
 A domain expert can perform this refinement relatively easily using the Microsoft Threat Modelling Tool, taking about 30 minutes for this example.
-The final curated list of threats for the IVI system is shown in Table 9 [8].
+The final curated list of threats for the IVI system is shown in Table 8 of our online appendix (see [10]).
 4.4 Step 4: Evaluate impact
@@
     • Privacy Impact: Deals with damages resulting from privacy violations.
-    Each of the four parameters is assigned one of four impact levels: None, Low, Medium, or High.
-For the purposes of this experience report, we used the impact level definitions from [12].
-Impact levels are typically assigned by automotive domain experts who have detailed knowledge of the automotive system, its vulnerabilities, and corresponding impacts on security objectives.
-In our running example, we assign impact levels based on our assessment of potential safety, operational, financial, and privacy consequences.
-This assessment is based on our understanding from the literature, and may differ from evaluations of experts in the field.
-    Next, numerical values are assigned to the impact levels.
-These values are chosen in a logarithmic scale (0, 1, 10, 100) to effectively capture the exponential increase in the impact levels.
-This helps in clearly differentiating between the levels of impact related to the security objectives, and ensures that high-impact threats are given appropriate priority.
+    Each of the four parameters is assigned an impact level: None, Low, Medium, or High, based on definitions from [11].
+These levels are typically determined by automotive domain experts with in-depth knowledge of the system and its vulnerabilities.
+For this report, we assigned impact levels based on our assessment of potential consequences, informed by literature, though these may differ from expert evaluations.
+    To quantify these levels, numerical values are assigned on a logarithmic scale (0, 1, 10, 100), reflecting the exponential increase in impact severity.
+This approach ensures clear differentiation between impact levels and prioritises high-impact threats effectively.
 Table 9 [10] categorises the impact levels and parameter values according to the security objectives.
   • Calculate impact score: After estimating the individual impact levels of the security objectives and their corresponding parameter values, the overall impact score is computed using the following equation:
-    $\small I=w_{s}i_{s}+w_{o}i_{o}+w_{f}i_{f}+w_{p}i_{p}$ (1)
+    $I_{\text{sum}}=w_{s}i_{s}+w_{o}i_{o}+w_{f}i_{f}+w_{p}i_{p}$ (1)
     Normalisation is necessary to ensure that the overall impact score is on a consistent scale, allowing for fair and meaningful comparisons across different systems and contexts.
@@
     The overall normalised impact score is computed using the following equation:
-    $I_{\text{norm}}=\frac{w_{s}i_{s}+w_{o}i_{o}+w_{f}i_{f}+w_{p}i_{p}}{100*(w_{s}+w_{o}+w_{f}+w_{p})}$ (2)
+    $I_{\text{nsum}}=\frac{w_{s}i_{s}+w_{o}i_{o}+w_{f}i_{f}+w_{p}i_{p}}{100*(w_{s}+w_{o}+w_{f}+w_{p})}$ (2)
     Weight values $w_{s}$, $w_{o}$, $w_{f}$, and $w_{p}$ are selected based on the specific needs of the use case being evaluated.
@@
 Equations 1 and 2 are simplified as follows:
-    $I=10i_{s}+i_{o}+i_{f}+i_{p}$ (3)
-    $\small I_{\text{norm}}=\frac{10i_{s}+i_{o}+i_{f}+i_{p}}{1300}$ (4)
-  • Assign impact rating: Each threat is assigned an impact rating based on the impact score $I_{\text{norm}}$ of each assessed threat, matched against thresholds tabulated in Table 2, as referenced from [12].
+    $I_{\text{sum}}=10i_{s}+i_{o}+i_{f}+i_{p}$ (3)
+    $I_{\text{nsum}}=\frac{10i_{s}+i_{o}+i_{f}+i_{p}}{1300}$ (4)
+  • Assign impact rating: Each threat is assigned an impact rating based on the impact score $I_{\text{nsum}}$, matched against thresholds tabulated in Table 1, as referenced from HEAVENS 2.0 [11].
+The unequal ranges of the impact scores emphasise the importance of distinguishing between high-severity and critical-severity impacts, particularly in safety-critical systems like automotive systems.
+Similarly, the smaller ranges for lower-severity impacts reflect the fact that these scenarios are less critical and require less granularity in their assessment.
     Table 1: Impact rating table
-    Impact Score Impact Rating
+    Impact Score ($I_{\text{nsum}}$) Impact Rating
     0.00 None
@@
 As such, the impact level for privacy is assessed as Medium and $i_{p}=10$.
-Using Equation 2, $I_{\text{norm}}\ =0.0077$, which corresponds to a Low impact rating.
+Using Equation 2, $I_{\text{nsum}}\ =0.0077$, which corresponds to a Low impact rating.
 Table 2: Impact rating example
@@
 9146 0 0 0 10 0.0077 Low
-4.5 Step 5: Model attacks
-In this step, we use an attack tree to model attacker goals and map out attack paths associated with threat and attack vectors identified earlier using the Microsoft Threat Modelling Tool.
+4.5 Step 5: Model attacks (cyber & physical)
+In this step, we use an attack tree to model attacker goals and map out attack paths associated with both physical and cyber threat and attack vectors.
 Since the vehicle driver is within the car trust boundary (Figure 2), we will assume the attacker personas and their goals as follows:
-  • Persona A: A car thief interested in stealing sensitive data from the In-Vehicle Infotainment system.
+  • Persona A: A car thief interested in stealing sensitive data from the IVI system.
   • Persona B: A rogue security researcher interested in manipulating vehicle functions that compromise the safety of the vehicle.
-The attack tree depicted in Figure 3 illustrates the potential attack paths these two personas could employ to achieve their respective goals.
-Figure 3: Attack tree for In-Vehicle Infotainment (IVI) system
+Figure 3: Attack tree for an In-Vehicle Infotainment (IVI) system, illustrating the combination of physical and cyber tactics to achieve the attacker’s goals
+The attack tree depicted in Figure 3 demonstrates how both personas could leverage a combination of physical and cyber tactics to achieve their objectives.
+One notable attack path included in the tree is the dirty road attack [14], where an attacker manipulates road signs (e.g.
+altering or obscuring traffic signals or directional signs) to mislead vehicle sensors.
+This physical manipulation could then be coupled with a cyber attack, such as exploiting a vulnerability in the IVI system to gain unauthorised access.
+By including both physical and cyber attack vectors, the attack tree underscores the importance of ACTISM’s holistic approach to security modelling, addressing not only cyber but also physical attack surfaces that could be exploited by malicious actors.
 4.6 Step 6: Evaluate attack feasibility
-Next, we use CVSS version 3.1 to analyse and compute attack feasibility.
+Next, we use CVSS (Common Vulnerability Scoring System) version 3.1 to analyse and compute attack feasibility.
 CVSS provides a standardised method for assessing the severity of vulnerabilities based on their impact and exploitability factors, as described below.
+We have chosen to use CVSS 3.1 instead of the more recent CVSS 4.0 because CVSS 3.1 is currently more widely adopted across our tools, processes, and industry standards.
+This ensures consistency in scoring and compatibility with existing vulnerability management systems.
   • Impact Metrics:
@@
 5 Updating the Security Model After a Vulnerability Disclosure
-ACTISM enhances security modelling by introducing a dynamic and iterative methodology, resulting in a threat-informed, evolving security model.
-This approach addresses evolving cybersecurity threats that necessitate the incorporation of new knowledge gained from security monitoring and threat intelligence throughout the system’s life cycle.
-We now apply this methodology to a real-world example using the Tesla Electric Vehicle (EV) jail-break vulnerability disclosure [14].
-Although this vulnerability is specific to the Tesla EV platform, ACTISM is designed to be applicable to all automotive systems (including non-EVs) and their relevant vulnerabilities.
-After gathering technical details from the vulnerability disclosure, we review the IVI system security model developed in the earlier section and update the CVSS metrics as necessary.
-For this vulnerability disclosure, there is no need to include additional assets in the security model or to revise the DFD (Figure 2) and attack tree (Figure 3).
-Although there is an opportunity to enhance the DFD and attack tree to include new attack vectors, such as cyber-physical attacks like keyless entry systems [15], we will not do so in this example.
-However, it is critical to reassess the CVSS metrics associated with Threat ID 9146, which correlates with the vulnerability disclosure.
+To demonstrate ACTISM’s dynamic and iterative methodology, we simulate a response to the Tesla Electric Vehicle (EV) jail-break vulnerability disclosure [15].
+While this vulnerability is specific to Tesla EVs, ACTISM is designed to be applicable to all automotive systems, including non-EVs, and their associated vulnerabilities.
+Using technical details from the disclosure, we review the IVI system security model developed earlier and update the CVSS metrics as needed.
+In this case, no additional assets, DFD updates (Figure 2), or attack tree revisions (Figure 3) are required.
+Although new attack vectors, such as cyber-physical attacks like dirty road attacks [14] or keyless entry systems [16], could be incorporated, we focus here on reassessing the CVSS metrics for Threat ID 9146, which aligns with the disclosed vulnerability.
+This process highlights ACTISM’s adaptability in responding to real-world threats.
 The CVSS metrics corresponding to Threat ID 9146 are updated as follows:
@@
 Lastly, the privacy impact is expected to increase from Low to High ($i_{p}=100$), as the privacy violation of multiple vehicle owners and drivers may lead to extensive media coverage and severe loss of market share.
-Using Equation 2 to compute the updated normalised impact score, $I_{\text{norm}}\ =0.1692$, which corresponds to a High impact rating (Table 8).
+Using Equation 2 to compute the updated normalised impact score, $I_{\text{nsum}}\ =0.1692$, which corresponds to a High impact rating (Table 7).
 Table 7: Updated impact rating following vulnerability disclosure
@@
 Ultimately, it empowers stakeholders to adopt additional precautions and improve the cybersecurity resilience of their vehicles throughout their life cycle.
-6 Future directions
-While ACTISM provides an approach to dynamically updating security models with both cyber and physical threats, it is not without its limitations.
-One of the primary challenges is how the accuracy of the model can be validated given the constantly evolving nature of threats and the potential for incomplete or inaccurate threat intelligence data.
-Moreover, while ACTISM’s adaptability to new threats is a strength, it can also pose a challenge in maintaining the models’ integrity and consistency over time.
-Despite these limitations, the potential future directions outlined below can address these challenges and enhance the robustness of ACTISM.
-Ongoing threat intelligence gathering, including real-time monitoring and tracking of vulnerability disclosures, is crucial for keeping the security model current.
-Establishing a comprehensive repository of vulnerability attributes, stored in a standardised format such as XML or JSON, can enable automation and streamline the updating process.
-Additionally, formalising security models using Domain-Specific Languages (DSLs) like Automation Markup Language (AutomationML) and Meta Attack Language (MAL) can provide clearer descriptions of threats and vulnerabilities.
-Incorporating adversary tactics from the MITRE ATT&CK knowledge base can further enhance these models by offering insights into real-world attack strategies.
-Moreover, leveraging machine learning approaches to prioritise vulnerabilities based on their impact, utilising historical data and threat intelligence, can significantly improve automotive cybersecurity.
-These potential future directions can ensure that automotive cybersecurity remains resilient and adaptive to emerging cybersecurity challenges.
+6 Practitioner Survey and Discussion
+To assess the strengths of the ACTISM framework and identify areas for improvement, particularly in addressing emerging threats and enhancing usability in real-world automotive systems, feedback was gathered through a survey of 16 industry experts in Singapore.
+These experts included Chief Information Security Officers (CISOs) and cybersecurity practitioners specialising in security modelling and risk management.
+Representing a diverse range of industries across both private and public sectors, they provided a broad perspective on the evaluation of ACTISM.
+All respondents rated ACTISM as relevant, with a mean rating of 4.5 out of 5 based on a Likert scale, highlighting its potential as a robust framework for automotive cybersecurity.
+Furthermore, practitioners strongly agreed that ACTISM effectively integrates cyber and physical threat modelling (mean rating of 4.7 out of 5) and accurately identifies and models cyber-physical threats (mean rating of 4.1 out of 5).
+However, some respondents highlighted limitations in ACTISM, particularly its reliance on continuous threat intelligence updates to maintain effectiveness.
+They also noted challenges in implementing and using ACTISM compared to other security modelling frameworks.
+For automotive manufacturers with limited access to real-time cybersecurity threat intelligence feeds, as well as the absence of automation in the model workflow, these limitations could result in inconsistencies, false positives, and reduced operational efficiency over time.
+To address these limitations and build on ACTISM’s strengths, we propose several future directions, many of which align with the recommendations provided by the experts.
+A key area for improvement is the need for adequate real-world or simulated testing to enhance the framework’s reliability.
+This would involve creating robust testing environments, including simulated attack scenarios, to evaluate ACTISM’s performance under realistic conditions.
+Such testing would also help validate the accuracy of the model, particularly given the constantly evolving nature of threats and the potential for incomplete or inaccurate threat intelligence data.
+Another recommendation from the experts is the integration of ACTISM with complementary frameworks such as the Exploit Prediction Scoring System (EPSS).
+Integrating EPSS, which predicts the likelihood of a vulnerability being exploited, alongside the Common Vulnerability Scoring System (CVSS), which measures the severity of a vulnerability, could enhance ACTISM’s ability to prioritise and respond to threats effectively.
+To further address the reliance on continuous threat intelligence updates, future work could focus on automating the model workflow.
+Establishing a comprehensive repository of vulnerability attributes, stored in a standardised format such as XML or JSON, would streamline the updating process and reduce the risk of inconsistencies.
+Additionally, formalising security models using Domain-Specific Languages (DSLs) like Automation Markup Language (AutomationML) and Meta Attack Language (MAL) could provide clearer descriptions of threats and vulnerabilities, improving the model’s usability and accuracy.
+Incorporating adversary tactics from the MITRE ATT&CK knowledge base would further enhance ACTISM by offering insights into real-world attack strategies.
+This integration would enable the framework to better anticipate and mitigate sophisticated threats.
+Moreover, leveraging machine learning approaches to prioritise vulnerabilities based on their impact, utilising historical data and threat intelligence, could significantly improve the framework’s predictive capabilities and overall effectiveness.
+These future directions will ensure that automotive cybersecurity remains resilient and capable of addressing emerging challenges in an increasingly complex threat landscape.
 7 Conclusion
 The cybersecurity landscape for automotive systems is rapidly evolving, with attackers increasingly targeting vulnerabilities that threaten data integrity, functionality, and safety.
-Traditional static security assessment methodologies, such as STRIDE, often fall short in addressing the dynamic and multi-faceted nature of these threats.
+The siloed application of traditional security assessment methodologies using STRIDE and other frameworks often limits their ability to address the evolving and multi-faceted (cyber & physical) nature of these threats.
 There is a clear knowledge gap in the ability to model and respond to security threats in a dynamic, iterative manner tailored specifically for automotive systems.
@@
 ACTISM effectively responds to evolving cyber-physical threats and attacker tactics, ensuring that security models remain robust and relevant throughout a system’s lifecycle.
-To illustrate the practical application of ACTISM, we developed a comprehensive security model for an IVI system.
-This process included identifying key assets, conducting thorough threat modelling, and performing risk assessments, ultimately resulting in a robust security framework.
-A real-world case study involving Tesla’s IVI system further demonstrated ACTISM’s ability to adapt to emerging threats and underscored its potential for broader applicability across various automotive technologies.
-In conclusion, ACTISM addresses the critical knowledge gap in automotive cybersecurity by providing a proactive, adaptable approach to security modelling.
-This approach not only tackles current challenges but also prepares for the future landscape of automotive technology.
+We applied ACTISM to Tesla’s In-Vehicle Infotainment system and showed that it can be practically used to model the dynamic nature of real-world cyber and physical threats in complex systems.
+Our experience report includes identifying key assets, conducting threat and attack modeling, and performing risk assessments for a complex system, in a dynamic, iterative manner.
+We also report the results of practitioners’ survey on the usefulness of ACTISM and its future directions.
+In conclusion, ACTISM bridges the critical gap in automotive cybersecurity by offering a proactive and adaptable approach to security modelling.
+This framework not only addresses current challenges but also prepares for the future, ensuring the resilience and safety of automotive systems in an ever-evolving threat landscape.